### v1

# Breaking the Dimensional Barrier: Dynamic Portfolio Choice with Parameter Uncertainty via Pontryagin Projection

Abstract
We study continuous-time portfolio choice in diffusion markets with parameter $\theta\in\Theta$ and uncertainty law $q(d\theta)$. Nature draws latent $\theta\sim q$ at time 0; the investor cannot observe it and must deploy a single $\theta$-blind feedback policy maximizing an ex–ante CRRA objective averaged over diffusion noise and $\theta$. Our methods access $q$ only by sampling and assume no parametric form. We extend Pontryagin–Guided Direct Policy Optimization (PG–DPO) by sampling $\theta$ inside the simulator and computing discrete-time gradients via backpropagation through time (BPTT), and we propose projected PG–DPO (P–PGDPO) that projects costate estimates to satisfy the $q$-aggregated Pontryagin first-order condition, yielding a deployable rule. We prove a BPTT–PMP correspondence uniform on compacts and a residual-based $\theta$-blind policy-gap bound under local stability with explicit discretization/Monte Carlo errors; experiments show projection-driven stability and accurate decision-time benchmark recovery in high dimensions.
1 Introduction
A central problem in quantitative finance is to allocate wealth dynamically across many risky assets in continuous time. In the classical Merton model, investment opportunities are described by a low-dimensional diffusion with known drift and volatility, and the investor solves a Hamilton–Jacobi–Bellman (HJB) equation to obtain closed-form optimal portfolios and value functions; see, for example, merton1969lifetime; merton1971optimum and the subsequent literature. In realistic markets, however, neither expected returns nor volatilities are known: they must be estimated from finite samples, often using many assets and predictors and under nontrivial model selection and regularization. Empirically, expected-return forecasting is fragile and return predictability is unstable across samples and over time, with many proposed predictors delivering limited out-of-sample gains (e.g., goyalwelch2008; campbellthompson2008; rapach2010forecast; danglhalling2012tvp; lettauvanNieuwerburgh2008shifts; pettenuzzo2014constraints). In such settings it is crucial to distinguish diffusion risk (Brownian noise conditional on parameters) from statistical parameter uncertainty (error in estimated coefficients). A long line of portfolio-choice work shows that return predictability, learning, and parameter uncertainty can induce substantial intertemporal hedging effects and more conservative allocations (e.g., kandelstambaugh1996predictability; barberis2000investor; campbell2002strategic; brandt2005portfolio; brennan1998investor; xia2001learning; maenhout2004robust).
We study continuous-time portfolio choice when market dynamics are known only up to an estimated parameter $\theta\in\Theta$, where the estimation pipeline produces a nondegenerate uncertainty law $q(d\theta)$ over $\Theta$. We treat $q$ as an input object that encapsulates all statistical information available at time 0: it may be derived from resampling approximations (e.g. bootstrap) (e.g., efron1979bootstrap; efron1994bootstrap), model averaging or Bayesian model uncertainty pipelines (e.g., pastor2000portfolio; avramov2002stock; cremers2002stock), approximate posteriors, or other procedures. Our goal is not to revisit inference, but to optimize decisions given this uncertainty description. Algorithmically, we interact with $q$ only through sampling $\theta\sim q$ inside the simulator; we do not assume closed-form densities, conjugate updates, or any particular parametric form. Concretely, we seek portfolio policies that maximize terminal CRRA utility ex–ante, averaging over both diffusion noise and parameter draws $\theta\sim q$. This formulation also supports offline diagnostics that quantify how recommended allocations vary across the statistically plausible models encoded in $q$.
A key modeling choice is that $\theta$ is latent: Nature draws a fixed but unobserved $\theta\sim q$ at time 0 (independent of the Brownian drivers) and keeps it fixed on $[0,T]$. While the investor knows $q$, she does not observe the realized $\theta$ and must therefore deploy a single $\theta$-blind policy. We restrict attention to Markov feedback policies of the form $\pi_{t}=\bar{\pi}(t,X_{t},Y_{t})$ that depend only on observable wealth $X_{t}$ and market factors $Y_{t}$, and we do not augment the state by a belief/posterior process. This fixed-$q$ commitment is intended as a decision-time benchmark: it targets a single deployable rule given an exogenous uncertainty law, cleanly decoupling how uncertainty is produced (any pipeline yielding $q$) from how decisions are optimized (our solver given $q$). We do not claim that belief-state control is conceptually inappropriate; rather, it defines a different (and typically far more demanding) problem than computing a single $\theta$-blind deployable feedback rule from a fixed uncertainty description (e.g., bensoussan1985optimal; pham2017dynamic). At the same time, fixed-$q$ optimization couples heterogeneous market models: gradient signals can vary substantially across $\theta$ draws and may partially cancel when learning a single global policy end-to-end.
The $\theta$-blind constraint also changes what a first-order optimality condition means. If $\theta$ were observable, Pontryagin’s Maximum Principle (PMP) yields a $\theta$-conditional criticality condition and an associated $\theta$-conditional full-information feedback map (infeasible under latent $\theta$). Under $\theta$-blind deployability, admissible perturbations are also $\theta$-blind. Taking the first variation of the ex–ante objective and using Fubini’s theorem shows that the correct necessary condition is $q$-aggregated: the expectation over $\theta\sim q$ of the Hamiltonian gradient $\partial_{\pi}H^{\mathrm{ctrl}}_{\theta}$ must vanish along the state process, in the standard stochastic maximum principle framework (e.g., yong1999stochastic; fleming2006controlled; pham2009continuous). Because $\partial_{\pi}H^{\mathrm{ctrl}}_{\theta}$ is affine in $\pi$ for our portfolio Hamiltonian, this aggregation yields a statewise linear system whose solution defines a deployable $\theta$-blind projected portfolio rule. Notably, the condition and resulting projection are agnostic to the internal construction of $q$ and depend only on its role as the ex–ante mixing law.
These features place the problem outside the practical reach of classical dynamic programming in the high-dimensional regime we target. In low-dimensional deterministic-parameter Markov models, DP/HJB is canonical; however, even with several factors it requires solving an HJB equation in the state $(t,X_{t},Y_{t})$, where grid-based PDE methods are quickly defeated by the curse of dimensionality (e.g., bellman1961adaptive; kushner2001numerical). Deep PDE surrogates such as PINNs (e.g., raissi2019physics; sirignano2018dgm) and deep BSDE methods (e.g., han2018solving; beck2019machine) alleviate the need for grids, but fully nonlinear portfolio HJBs with many assets and factors remain numerically delicate, especially when accurate mixed derivatives are required. If one further models parameter uncertainty via belief-state augmentation, the state becomes a posterior measure and the control problem becomes infinite-dimensional (e.g., bensoussan1985optimal; pham2017dynamic).
Our approach is simulation-based and builds on Pontryagin–Guided Direct Policy Optimization (PG–DPO) (huh2025breaking; huh2025constraint). PG–DPO parameterizes a $\theta$-blind feedback policy via a neural network, simulates trajectories of the controlled SDE, and employs backpropagation through time (BPTT) to compute exact gradients of terminal utility. Crucially, intermediate pathwise sensitivities computed by BPTT coincide with the stochastic costates (adjoints) in PMP, mirroring the classical duality between backpropagation and adjoint methods (see lecun1988optimal; yong1999stochastic). In the latent-parameter setting, we approximate the ex–ante objective by sampling $\theta\sim q$ inside the simulator and fixing it along each trajectory, while the policy depends only on observable states. To stabilize learning under heterogeneous $\theta$ draws, we extend the projected variant, P–PGDPO, to latent $\theta$: after a warm-up phase that stabilizes costate estimates, we project Monte Carlo Pontryagin objects onto the $q$-aggregated Pontryagin first-order condition. This reconstruction yields a robust deployable $\theta$-blind rule obtained from the $q$-aggregated criticality, and can be amortized into a fast-to-evaluate policy.
In high-dimensional scaling experiments under static Gaussian drift uncertainty, the two-stage projected pipeline substantially improves decision-time accuracy relative to end-to-end learning, with clear stabilization effects in aligned regimes. In misaligned regimes, projection gains diminish with dimension; diagnostics indicate that deterioration is driven primarily by growth of aggregated first-order residuals and curvature mismatch rather than by catastrophic numerical inversion. In factor-driven markets with mean-reverting investment opportunities where return–factor correlation induces intertemporal hedging demand, the projected pipeline recovers the analytic decision-time benchmark under the same $\theta$-blind deployability restriction, while a model-free PPO baseline remains far from the reference in the regimes we test.
Our main theoretical guarantee is a residual-based ex–ante $\theta$-blind policy-gap bound for the deployable fixed-$q$ commitment problem: under mild slab-wise local stability conditions for the $q$-aggregated projection map, a small warm-up aggregated first-order residual implies that the projected policy is close (in $L^{2}(\mu)$) to a locally optimal interior deployable $\theta$-blind policy, up to discretization and Monte Carlo error.
Our contributions are threefold. (i) We formulate a latent-parameter, fixed-$q$ ex–ante CRRA portfolio problem under a deployable $\theta$-blind Markov feedback restriction and derive the corresponding $q$-aggregated Pontryagin first-order condition, emphasizing an inference-agnostic interface where uncertainty enters only through an exogenous mixing law $q(d\theta)$. (ii) We extend PG–DPO to this setting by sampling $\theta$ only inside the simulator and using BPTT to compute exact discrete-time gradients and pathwise sensitivities, and we establish a conditional BPTT–PMP correspondence uniform over $\theta$ on compact subsets of $\Theta$. (iii) We develop uncertainty-aware P–PGDPO that projects Monte Carlo costate estimates to produce a deployable $q$-aggregated $\theta$-blind rule, together with a residual-based ex–ante $\theta$-blind policy-gap bound and empirical evidence of two-time-scale stabilization and stability gains from projection.
The remainder of the paper is organized as follows. Section 2 formulates the fixed-$q$ ex–ante portfolio problem under a latent parameter and a deployable $\theta$-blind Markov feedback restriction, and derives the $\theta$-conditional versus $q$-aggregated Pontryagin first-order conditions together with Gaussian decision-time reference models. Section 3 develops PG–DPO and uncertainty-aware P–PGDPO for the latent-$\theta$ setting, establishes the conditional BPTT–PMP correspondence, and proves a residual-based ex–ante $\theta$-blind policy-gap bound under local stability of the aggregated projection map. Section 4 reports high-dimensional scaling experiments under static Gaussian drift uncertainty, and Section 5 studies hedging-demand recovery in factor-driven markets with mean-reverting investment opportunities. Technical proofs and implementation details are collected in the appendix.
2 Dynamic Portfolio Choice in Estimated Diffusion Markets with Latent Parameter Uncertainty
In this section we formulate a continuous-time dynamic portfolio choice problem with CRRA preferences in a diffusion market whose coefficients are estimated from data and are therefore statistically uncertain. Rather than committing to a particular estimation architecture, we treat the market as belonging to a (possibly high-dimensional) parameterized family indexed by $\theta\in\Theta$, and we represent the uncertainty in the estimated parameter by an exogenously given probability law $q(d\theta)$ over $\Theta$.
  • Nature draws a fixed but unobserved $\theta\sim q$ at time 0 and keeps it constant on $[0,T]$.
  • The investor knows $q$ but does not observe the realized $\theta$ and must deploy a single $\theta$-blind portfolio policy.
  • Performance is evaluated ex–ante by averaging terminal utility over both diffusion noise and $\theta\sim q$.
  • We restrict to $\theta$-blind Markov feedback rules $\pi_{t}=\bar{\pi}(t,X_{t},Y_{t})$ and do not augment the state by a belief/posterior process.
This fixed-$q$, $\theta$-blind formulation is intentionally algorithm-facing: we view the estimation procedure that produced $q(d\theta)$ as exogenous, and our goal is to compute stable, scalable portfolio rules given this uncertainty description. It is also a commitment model: the investor commits at time 0 to a single feedback map and does not update $q$ during trading. As a result, one must distinguish between (i) $\theta$-conditional (full-information) optimality conditions and objects that would be available if $\theta$ were observable (infeasible under latent $\theta$), and (ii) $q$-aggregated conditions that characterize optimality within the $\theta$-blind admissible class. Throughout, $\theta$-conditional objects are used only for offline diagnostics (e.g., heterogeneity inspection and infeasible upper bounds), whereas our algorithms target a single deployable $\theta$-blind rule; see Section 2.2.
2.1 Model and ex–ante objective in estimated diffusion markets
We work on a filtered probability space $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\in[0,T]},\mathbb{P})$ supporting Brownian motions of appropriate dimension. Time is continuous and runs over a fixed finite horizon $[0,T]$.
Deterministic-parameter reference (classical CRRA Merton).
There is one risk-free asset (money market account) with price process $B$ satisfying
$\frac{dB_{t}}{B_{t}}=r\,dt,\hskip 18.49988ptB_{0}=1,$ (1)
where $r\in\mathbb{R}$ is a constant short rate. In the classical Merton model, the $d$ risky assets have prices $S_{t}=(S_{t}^{1},\dots,S_{t}^{d})^{\top}$ solving
$\frac{dS_{t}}{S_{t}}:=\bigg(\frac{dS_{t}^{1}}{S_{t}^{1}},\dots,\frac{dS_{t}^{d}}{S_{t}^{d}}\bigg)^{\top}=r\,\mathbf{1}\,dt+\mu\,dt+\Sigma^{1/2}dW_{t},\hskip 18.49988ptS_{0}\in(0,\infty)^{d},$ (2)
with constant excess returns $\mu\in\mathbb{R}^{d}$, volatility matrix $\Sigma^{1/2}\in\mathbb{R}^{d\times d}$, and a $d$-dimensional Brownian motion $W$. An investor with CRRA utility $U(x)=x^{1-\gamma}/(1-\gamma)$, $\gamma>0,\gamma\neq 1$, chooses a progressively measurable portfolio fraction $\pi_{t}\in\mathbb{R}^{d}$; the wealth process satisfies
$\frac{dX_{t}^{\pi}}{X_{t}^{\pi}}=\big(r+\pi_{t}^{\top}\lambda\big)dt+\pi_{t}^{\top}\Sigma^{1/2}dW_{t},\hskip 18.49988ptX_{0}^{\pi}=x>0,$ (3)
where $\lambda:=\mu$ is the vector of risk premia. In this benchmark setting the optimal policy is constant:
$\pi^{\star}=\frac{1}{\gamma}\,\Sigma^{-1}\lambda,$ (4)
and the corresponding value function is given explicitly by
$V^{\mathrm{Merton}}(t,x;\lambda)=\frac{x^{1-\gamma}}{1-\gamma}\,\exp\Big\{(1-\gamma)\Big(r+\tfrac{1}{2\gamma}\lambda^{\top}\Sigma^{-1}\lambda\Big)(T-t)\Big\},$ (5)
see, for example, merton1969lifetime; merton1971optimum. We use this constant-coefficient model only as a deterministic-parameter reference.
Estimated diffusion market family (conditional on a latent parameter).
In our main formulation, drift and volatility are not assumed known. Instead, we consider a general multi-asset, multi-factor diffusion family indexed by $\theta\in\Theta\subset\mathbb{R}^{k}$, where $\theta$ represents the (possibly high-dimensional) parameter produced by an estimation procedure. Conditional on $\theta$, the $d$ risky assets and an $m$-dimensional factor process $Y_{t}$ evolve as
$\frac{dS_{t}}{S_{t}}$ $=r\,\mathbf{1}\,dt+b\big(Y_{t},\theta\big)\,dt+\sigma\big(Y_{t},\theta\big)\,dW_{t},\hskip 18.49988ptS_{0}\in(0,\infty)^{d},$ (6)
$dY_{t}$ $=a\big(Y_{t},\theta\big)\,dt+\beta\big(Y_{t},\theta\big)\,dW_{t}^{Y},\hskip 18.49988ptY_{0}=y\in\mathbb{R}^{m},$ (7)
where $W$ and $W^{Y}$ are Brownian motions (possibly of different dimension) that may be instantaneously correlated. We write the instantaneous covariance and return–factor cross-covariance as
$\Sigma(y,\theta):=\sigma(y,\theta)\sigma(y,\theta)^{\top},\hskip 18.49988pt\Sigma_{SY}(y,\theta):=\sigma(y,\theta)\,\rho\,\beta(y,\theta)^{\top},$ (8)
where $\rho$ is defined by $d\langle W,W^{Y}\rangle_{t}=\rho\,dt$. Thus $\Sigma(y,\theta)\in\mathbb{R}^{d\times d}$ and $\Sigma_{SY}(y,\theta)\in\mathbb{R}^{d\times m}$.
Uncertainty law $q(d\theta)$ and information structure.
The parameter $\theta$ is estimated from finite samples and is uncertain. We summarize this uncertainty by a probability distribution
$q(d\theta).$ (9)
We deliberately do not tie $q$ to any specific inference paradigm. Concretely, $q$ may represent an empirical/sampling distribution produced by resampling procedures such as the bootstrap (efron1979bootstrap; efron1994bootstrap), a distribution induced by model averaging or sub-sample aggregation procedures such as bagging (breiman1996bagging), an approximate Bayesian posterior (when a prior and likelihood/criterion are specified), or an asymptotic normal (or sandwich) approximation in parametric or semiparametric estimation. For our purposes, $q$ is an input object describing statistically plausible market parameters.
Remark 1 (Latent parameter, observability, and admissible controls).
We interpret $\theta$ as a latent (unobserved) market parameter: Nature draws an $\mathcal{F}_{0}$-measurable random variable $\theta\sim q$ at time 0 (independent of the Brownian drivers) and keeps it fixed over $[0,T]$. The investor knows $q$ but does not observe the realized $\theta$, so deployable portfolio rules cannot take $\theta$ as an input.
We consider the observable market filtration
$\mathcal{F}_{t}^{\mathrm{obs}}:=\sigma\{(S_{s},Y_{s}):0\leq s\leq t\},\hskip 18.49988pt0\leq t\leq T,$ (10)
where $\sigma\{\cdot\}$ denotes the $\sigma$-field generated by the observed asset and factor paths (with the usual augmentation). Admissible portfolio processes are required to be progressively measurable with respect to $(\mathcal{F}_{t}^{\mathrm{obs}})$.
Throughout the paper we restrict attention to the Markov feedback subclass
$\mathcal{A}^{\mathrm{fb}}:=\Big\{\pi\in\mathcal{A}^{\mathrm{obs}}:\ \exists\,\bar{\pi}:[0,T]\times(0,\infty)\times\mathbb{R}^{m}\to\mathbb{R}^{d}\ \text{s.t.}\ \pi_{t}=\bar{\pi}(t,X_{t},Y_{t})\Big\},$ (11)
where $\mathcal{A}^{\mathrm{obs}}$ is defined below. This restriction reflects a fixed-$q$ commitment model: the investor uses historical data to form $q$ prior to trading and does not perform online filtering/belief-state updates during $[0,T]$.
Whenever we display $\theta$-conditional (full-information) controls or sensitivity objects, they are computed under frozen-$\theta$ simulations and are used only for offline diagnostics; the deployed policy class and the learned policy remain $\theta$-blind.
Wealth dynamics and admissibility (given $\theta$).
For any fixed $\theta$, the corresponding wealth dynamics under a portfolio process $\pi_{t}(\omega)\in\mathbb{R}^{d}$ adapted to $\mathcal{F}_{t}^{\mathrm{obs}}$ are
$\frac{dX_{t}^{\pi}}{X_{t}^{\pi}}=\Big(r+\pi_{t}^{\top}b(Y_{t},\theta)\Big)dt+\pi_{t}^{\top}\sigma(Y_{t},\theta)\,dW_{t},$ (12)
and we denote by $\mathcal{A}^{\mathrm{obs}}$ the set of progressively measurable portfolio processes adapted to $(\mathcal{F}_{t}^{\mathrm{obs}})$ for which (12) admits a (strictly) positive wealth solution. In the Markovian feedback case $\pi\in\mathcal{A}^{\mathrm{fb}}$ one may think of $\pi_{t}=\bar{\pi}(t,X_{t},Y_{t})$.
Ex–ante objective under latent $\theta$ (and simulator viewpoint).
The investor evaluates policies under an ex–ante objective that averages over both diffusion noise for fixed $\theta$ and the parametric uncertainty encoded by (9):
$J(\pi):=\mathbb{E}_{\theta\sim q}\bigg[\mathbb{E}\big[U(X_{T}^{\pi})\,\big|\,\theta\big]\bigg]=\int_{\Theta}\mathbb{E}\big[U(X_{T}^{\pi})\,\big|\,\theta\big]\,q(d\theta).$ (13)
The corresponding optimization problem (under our feedback restriction) is
$\sup_{\pi\in\mathcal{A}^{\mathrm{fb}}}J(\pi).$ (14)
Whenever it exists, we denote by
$\pi^{\star,\mathrm{blind}}\in\arg\max_{\pi\in\mathcal{A}^{\mathrm{fb}}}J(\pi)$
an optimal $\theta$-blind feedback for the fixed-$q$ commitment problem (14). For each fixed $\theta$, we also write $\pi^{\star,\theta}$ for the (infeasible) $\theta$-conditional full-information optimal control that would be available if $\theta$ were observed.
The $\theta$-blind constraint makes (14) strictly harder than solving a separate control problem for each fixed $\theta$, since the latter yields a $\theta$-indexed full-information family. Ex–ante averaging in (13) can also create gradient cancellation across heterogeneous parameter draws when one attempts to learn a single global policy end-to-end. While an $\mathcal{F}_{t}^{\mathrm{obs}}$-adapted policy could, in principle, filter $\theta$ online and solve a belief-state control problem (see, e.g., bensoussan1985optimal; pham2017dynamic), we do not pursue that formulation here.
Approximating the outer expectation in (13) amounts to sampling $\theta\sim q$ inside the simulator (once per trajectory or once per update), running (6)–(7) under that frozen draw, and updating a $\theta$-blind feedback policy to perform well on average over such draws. This is the setting targeted by the simulation-based PG–DPO and P–PGDPO methods developed in Section 3.
2.2 Pontryagin optimality under latent parameters: full-information vs. aggregated conditions
This subsection records the Hamiltonian structure underlying our projection step and clarifies what “Pontryagin first-order conditions” mean when the market parameter $\theta$ is latent and admissible controls are $\theta$-blind. In particular, we distinguish between (i) $\theta$-conditional (full-information) criticality conditions that would apply if $\theta$ were observable (and are therefore infeasible under latent $\theta$), and (ii) $q$-aggregated criticality conditions that characterize stationarity within the $\theta$-blind admissible class for the fixed-$q$ ex–ante objective. Our discussion follows standard stochastic control/PMP arguments for diffusion control (e.g. yong1999stochastic; fleming2006controlled; pham2009continuous). We also comment on the relationship to partial-information (belief-state) PMP, but we do not develop that formulation here.
A $\theta$-conditional (full-information) Hamiltonian and first-order condition (infeasible under latent $\theta$).
Fix $\theta\in\Theta$ and suppose, for the moment, that $\theta$ were observable to the controller. In Markovian settings with sufficient smoothness, the $\theta$-conditional value function $V^{\star,\theta}(t,x,y)$ satisfies an HJB equation whose control Hamiltonian (the part depending on $\pi$) can be written explicitly using (8):
$\mathcal{H}_{\theta}^{\mathrm{ctrl}}(t,x,y,\pi;\,V_{x},V_{xx},V_{xy}):=x\,\pi^{\top}b(y,\theta)\,V_{x}+\frac{1}{2}x^{2}\,\pi^{\top}\Sigma(y,\theta)\,\pi\,V_{xx}+x\,\pi^{\top}\Sigma_{SY}(y,\theta)\,V_{xy},$ (15)
where $V_{x},V_{xx}$ are evaluated at $(t,x,y)$ and $V_{xy}(t,x,y)\in\mathbb{R}^{m}$. The last term in (15) is the return–factor hedging term induced by $d\langle W,W^{Y}\rangle\neq 0$.
The pointwise first-order condition for an interior optimizer is
$\partial_{\pi}\mathcal{H}_{\theta}^{\mathrm{ctrl}}=x\,V_{x}^{\star,\theta}\,b(y,\theta)+x^{2}\,V_{xx}^{\star,\theta}\,\Sigma(y,\theta)\,\pi+x\,\Sigma_{SY}(y,\theta)\,V_{xy}^{\star,\theta}\;=\;0.$ (16)
Assuming $\Sigma(y,\theta)$ is invertible and $V_{xx}^{\star,\theta}<0$, this yields the closed-form $\theta$-conditional full-information portfolio rule
$\pi^{\star,\theta}(t,x,y)=-\,\frac{1}{x\,V_{xx}^{\star,\theta}(t,x,y)}\,\Sigma(y,\theta)^{-1}\Big(V_{x}^{\star,\theta}(t,x,y)\,b(y,\theta)+\Sigma_{SY}(y,\theta)\,V_{xy}^{\star,\theta}(t,x,y)\Big).$ (17)
This $\theta$-indexed rule is not deployable under latent parameters; we record it only as a full-information benchmark and diagnostic reference. In our setting, deployable policies never take the realized $\theta$ as an input; $\theta$ is accessed only through sampling inside the simulator when approximating $q$-expectations.
$q$-aggregated Pontryagin condition for the $\theta$-blind ex–ante problem (Markov feedback).
We now return to the actual setting: $\theta$ is latent, policies are $\theta$-blind, and we restrict attention to the Markov feedback class $\mathcal{A}^{\mathrm{fb}}$ (Remark 1). Under this restriction we neither perform online filtering of $\theta$ nor replace $q$ by a time-varying posterior distribution. Accordingly, the relevant Pontryagin condition is not the $\theta$-conditional criticality (16) enforced pointwise in $\theta$, but rather a necessary condition for optimality within the $\theta$-blind admissible class for the fixed-$q$ objective (13).
To see why ex–ante aggregation enters the first-order condition, take any $\theta$-blind admissible perturbation $h=\{h_{t}\}_{t\in[0,T]}$ that is progressively measurable with respect to the observation filtration $(\mathcal{F}_{t}^{\mathrm{obs}})$ and square-integrable, and define $\pi^{\varepsilon}:=\pi+\varepsilon h$ for small $\varepsilon$. For each fixed $\theta$, the stochastic maximum principle yields the first-variation identity
$\left.\frac{d}{d\varepsilon}J^{\theta}(\pi^{\varepsilon})\right|_{\varepsilon=0}=\mathbb{E}\!\left[\int_{0}^{T}\partial_{\pi}\mathcal{H}^{\mathrm{ctrl}}_{\theta}\big(t,X_{t},Y_{t},\pi_{t};\,p_{t}^{\theta},p_{x,t}^{\theta},p_{y,t}^{\theta}\big)^{\top}h_{t}\,dt\,\Big|\,\theta\right],$ (18)
where $\big(p_{t}^{\theta},p_{x,t}^{\theta},p_{y,t}^{\theta}\big)$ denotes the $\theta$-conditional Pontryagin sensitivity objects associated with the fixed policy $\pi$ in the frozen-$\theta$ market. Because both $\pi$ and $h$ are $\theta$-blind, taking the outer expectation over $\theta\sim q$ and using Fubini’s theorem gives
$\left.\frac{d}{d\varepsilon}J(\pi^{\varepsilon})\right|_{\varepsilon=0}=\mathbb{E}\!\left[\int_{0}^{T}\mathbb{E}_{\theta\sim q}\!\Big[\partial_{\pi}\mathcal{H}^{\mathrm{ctrl}}_{\theta}\big(t,X_{t},Y_{t},\pi_{t};\,p_{t}^{\theta},p_{x,t}^{\theta},p_{y,t}^{\theta}\big)\Big]^{\top}h_{t}\,dt\right].$ (19)
Hence, for an interior $\theta$-blind optimum $\pi^{\star,\mathrm{blind}}$, the first variation must vanish for all such perturbations $h$, which implies the aggregated first-order condition
$\mathbb{E}_{\theta\sim q}\!\Big[\partial_{\pi}\mathcal{H}^{\mathrm{ctrl}}_{\theta}\big(t,X_{t},Y_{t},\pi_{t};\,p_{t}^{\theta},p_{x,t}^{\theta},p_{y,t}^{\theta}\big)\Big]=0,\hskip 18.49988pt\text{a.s.\ for a.e.\ }t\in[0,T].$ (20)
Equation (20) is the correct necessary condition for the ex–ante problem under the $\theta$-blind constraint. In particular, it is generally distinct from imposing (16) for each $\theta$ separately, because $\theta$-conditional criticality cannot be enforced by a single deployable $\theta$-blind policy.
To operationalize (20) in the Markov feedback class, fix a feedback policy $\pi\in\mathcal{A}^{\mathrm{fb}}$ and, for each frozen $\theta$, consider the corresponding $\theta$-conditional Pontryagin sensitivity objects $\big(p_{t}^{\theta},p_{x,t}^{\theta},p_{y,t}^{\theta}\big)$ along the induced state process. In smooth Markov regimes these coincide with spatial derivatives of a decoupling field and, in particular, reduce to $(V_{x},V_{xx},V_{xy})$ in the full-information setting; in our algorithms we estimate them pathwise by automatic differentiation (see Section 3).
For the portfolio Hamiltonian (15), $\partial_{\pi}\mathcal{H}_{\theta}^{\mathrm{ctrl}}$ is affine in $\pi$. This motivates defining the $\theta$-conditional “projection inputs”
$A_{t}^{\theta}(t,x,y)$ $:=x\,p_{x,t}^{\theta}(t,x,y)\,\Sigma(y,\theta)\in\mathbb{R}^{d\times d},$ (21)
$G_{t}^{\theta}(t,x,y)$ $:=p_{t}^{\theta}(t,x,y)\,b(y,\theta)+\Sigma_{SY}(y,\theta)\,p_{y,t}^{\theta}(t,x,y)\in\mathbb{R}^{d},$ (22)
and their $q$-aggregated counterparts
$A_{t}(t,x,y):=\mathbb{E}_{\theta\sim q}\!\big[A_{t}^{\theta}(t,x,y)\big],\hskip 18.49988ptG_{t}(t,x,y):=\mathbb{E}_{\theta\sim q}\!\big[G_{t}^{\theta}(t,x,y)\big].$ (23)
These objects summarize how the latent parameter affects the first-order stationarity condition through the $\theta$-conditional sensitivities.
Theorem 1 ($q$-aggregated first-order condition under latent $\theta$ (deployable $\theta$-blind stationarity)).
Consider the fixed-$q$ ex–ante objective (13) over the $\theta$-blind Markov feedback class $\mathcal{A}^{\mathrm{fb}}$. Assume standard smoothness/integrability conditions ensuring validity of first variations within $\mathcal{A}^{\mathrm{fb}}$ and existence of the associated $\theta$-conditional Pontryagin objects. If $\pi^{\star,\mathrm{blind}}$ is a locally optimal interior policy in $\mathcal{A}^{\mathrm{fb}}$, then (20) holds. Moreover, in the portfolio setting (15), the aggregated stationarity is equivalent to the statewise linear system
$A_{t}(t,x,y)\,\pi^{\star,\mathrm{blind}}(t,x,y)=-\,G_{t}(t,x,y),\hskip 18.49988pt(t,x,y)\in[0,T]\times(0,\infty)\times\mathbb{R}^{m},$ (24)
(where $A_{t},G_{t}$ are defined by (23) using the $\theta$-conditional Pontryagin objects generated by $\pi^{\star,\mathrm{blind}}$). Whenever $A_{t}(t,x,y)$ is invertible on the working domain, (24) is equivalently expressed as the projected feedback rule
$\pi^{\mathrm{agg}}(t,x,y)=-\,A_{t}(t,x,y)^{-1}\,G_{t}(t,x,y).$ (25)
Proof sketch.
The conditional first-variation identity (18) is standard for diffusion control under a fixed parameter $\theta$ (e.g. yong1999stochastic; fleming2006controlled; pham2009continuous). Taking the outer expectation over $\theta\sim q$ yields (19). Since $h$ is an arbitrary $\theta$-blind admissible perturbation, vanishing of the first variation at an interior optimum implies (20). For the quadratic portfolio Hamiltonian (15), substituting the explicit expression for $\partial_{\pi}\mathcal{H}^{\mathrm{ctrl}}_{\theta}$ and introducing (21)–(23) yields the linear system (24) and the projected form (25) whenever $A_{t}$ is invertible. ∎
Note that $\pi^{\mathrm{agg}}$ is generally not equal to the naive average $\mathbb{E}_{\theta\sim q}[\pi^{\star,\theta}(t,x,y)]$ of $\theta$-conditional full-information controls, reflecting the noncommutativity between averaging over $\theta$ and solving a first-order condition. In particular, even if one could compute $\pi^{\star,\theta}$ for each $\theta$, averaging these infeasible oracles does not, in general, enforce the deployable $q$-aggregated stationarity (20).
Remark 2 (Relation to belief-state/learning formulations).
If one allows history-dependent policies that explicitly infer $\theta$ from observed returns, a principled partial-information formulation introduces a time-varying posterior/belief state $q_{t}(\cdot)=\mathbb{P}(\theta\in\cdot\mid\mathcal{F}_{t}^{\mathrm{obs}})$. In such belief-state problems, the corresponding PMP/Hamiltonian criticality condition is expressed in terms of conditional expectations under $q_{t}$ (or, equivalently, conditional on $\mathcal{F}_{t}^{\mathrm{obs}}$); see, e.g., haussmann1987maximum; li1995general; baghery2007maximum. We do not pursue that learning/belief-state route here. Our algorithms and theory target the fixed-$q$, $q$-aggregated projection (25) under the $\theta$-blind Markov feedback restriction (52).
2.3 Gaussian references at a fixed decision time
This subsection collects Gaussian benchmarks that isolate decision-time statistical uncertainty and yield closed-form reference allocations. We fix a calendar decision time $t_{0}$ at which an external estimation pipeline outputs an uncertainty law $q_{t_{0}}(d\theta)$ for a risk-premium parameter, and we treat this law as an $\mathcal{F}_{t_{0}}$-measurable input for portfolio choice over the remaining horizon. This interface accommodates both Bayesian posterior/prior-like uncertainty descriptions (e.g., barberis2000investor; pastor2000portfolio) and frequentist sampling/resampling laws conditional on $\mathcal{F}_{t_{0}}$ (e.g., bootstrap or bagging) (e.g., efron1979bootstrap; efron1994bootstrap; breiman1996bagging). Throughout this subsection we work conditionally on $\mathcal{F}_{t_{0}}$, suppress conditioning by writing $q(d\theta)$, and shift the trading clock so that the decision time becomes 0 and the remaining horizon is $T$. These references are used as analytic targets and sanity checks for our numerical sections (Sections 4 and 5), rather than as characterizations of the unrestricted optimum of (14) over the full feedback class.
We present two decision-time references. Section 2.3.1 considers static drift uncertainty: a latent premium is drawn from $q$ once at time 0 and kept fixed on $[0,T]$, providing the controlled benchmark used in the high-dimensional scaling/geometry experiments of Section 4. Section 2.3.2 considers a mean-reverting (OU) premium with Gaussian initial uncertainty, which induces a horizon-dependent Gaussian law for the time-averaged premium and yields a tractable closed-form reference used in the hedging-demand recovery study of Section 5. For completeness, an online linear–Gaussian illustration that produces a time-varying uncertainty law $q_{t}$ via Kalman–Bucy filtering is deferred to Appendix A; it is included only to motivate a plug-in (receding-horizon) decision-time workflow in which $q_{t}$ is treated as an externally updated input at each decision time, rather than solving the fully optimal belief-state control problem (e.g., bensoussan1985optimal; pham2017dynamic).
2.3.1 Static Gaussian drift uncertainty
We start from a time-homogeneous Gaussian benchmark in which the (vector) risk premium is an unobserved static parameter drawn at the decision time. The agent commits to a single $\theta$-blind policy, and all ex–ante uncertainty is summarized by the decision-time law $q$.
Market model (static latent drift).
Let $d$ risky assets satisfy
$\frac{dS_{t}}{S_{t}}=r\,\mathbf{1}\,dt+\theta\,dt+\Sigma^{1/2}dW_{t},\hskip 18.49988ptS_{0}\in(0,\infty)^{d},$ (26)
where $\Sigma\in\mathbb{R}^{d\times d}$ is symmetric positive definite and the latent excess-return vector is drawn at time 0 as
$\theta\sim q(d\theta).$ (27)
A $\theta$-blind portfolio fraction process $\pi_{t}\in\mathbb{R}^{d}$ generates wealth
$\frac{dX_{t}^{\pi}}{X_{t}^{\pi}}=\big(r+\pi_{t}^{\top}\theta\big)\,dt+\pi_{t}^{\top}\Sigma^{1/2}\,dW_{t},\hskip 18.49988ptX_{0}^{\pi}=x>0,$ (28)
and we evaluate the ex–ante objective
$J(\pi):=\mathbb{E}_{\theta\sim q}\Big[\mathbb{E}\big[U(X_{T}^{\pi})\mid\theta\big]\Big].$ (29)
For reference, under full information and CRRA utility $U(x)=x^{1-\gamma}/(1-\gamma)$ ($\gamma>0,\gamma\neq 1$), the oracle Merton rule is $\pi^{\star}(\theta)=\frac{1}{\gamma}\Sigma^{-1}\theta$ (merton1969lifetime; merton1971optimum), which is infeasible here because $\theta$ is latent.
Analytic $q$-references via constant portfolios. To obtain a transparent closed-form benchmark that depends only on the decision-time law $q$, we temporarily restrict attention to constant portfolio fractions
$\pi_{t}\equiv\pi\in\mathbb{R}^{d}.$ (30)
This restriction is used solely to define an analytic $q$-reference; it is not imposed on the learning problem.
Log utility ($\gamma=1$).
Let $m_{\theta}:=\mathbb{E}_{\theta\sim q}[\theta]$. For constant $\pi$, the objective depends on $q$ only through $m_{\theta}$, and the optimal constant portfolio is
$\pi_{q,\log}^{\mathrm{const}}(T)=\Sigma^{-1}m_{\theta},$ (31)
which is independent of $T$ in this static benchmark. In the one-asset case ($d=1$, $\Sigma=\sigma^{2}$),
$\pi_{q,\log}^{\mathrm{const}}(T)=\frac{m_{\theta}}{\sigma^{2}}.$ (32)
CRRA ($\gamma\neq 1$): tilted optimality and Gaussian shrinkage.
For $\gamma\neq 1$ and constant $\pi$, conditional on $\theta$ the terminal wealth is lognormal, and
$J(\pi)=\frac{x^{1-\gamma}}{1-\gamma}\,\exp\Big\{(1-\gamma)rT-\tfrac{1}{2}\gamma(1-\gamma)T\,\pi^{\top}\Sigma\pi\Big\}\,M_{q}\big((1-\gamma)T\pi\big),$ (33)
where $M_{q}(u):=\mathbb{E}_{\theta\sim q}[\exp(u^{\top}\theta)]$ is the moment generating function of $q$. Any interior optimizer $\pi_{q,\gamma}^{\mathrm{const}}(T)$ satisfies the tilted first-order condition
$\gamma\,\Sigma\,\pi_{q,\gamma}^{\mathrm{const}}(T)=\nabla_{u}\log M_{q}(u)\Big|_{u=(1-\gamma)T\,\pi_{q,\gamma}^{\mathrm{const}}(T)}.$ (34)
If $q$ is Gaussian,
$\theta\sim\mathcal{N}(m_{\theta},P),\hskip 18.49988ptP\succeq 0,$ (35)
then $\nabla_{u}\log M_{q}(u)=m_{\theta}+Pu$ and the reference reduces to the linear system
$\big(\gamma\Sigma-(1-\gamma)T\,P\big)\,\pi_{q,\gamma}^{\mathrm{const}}(T)=m_{\theta},$ (36)
hence
$\pi_{q,\gamma}^{\mathrm{const}}(T)=\big(\gamma\Sigma-(1-\gamma)T\,P\big)^{-1}m_{\theta}.$ (37)
For $\gamma>1$, this takes the familiar shrinkage form
$\pi_{q,\gamma}^{\mathrm{const}}(T)=\big(\gamma\Sigma+(\gamma-1)T\,P\big)^{-1}m_{\theta},\hskip 18.49988pt(\gamma>1),$ (38)
and in one dimension ($d=1$, $\Sigma=\sigma^{2}$, $P=p$),
$\pi_{q,\gamma}^{\mathrm{const}}(T)=\frac{m_{\theta}}{\gamma\sigma^{2}+(\gamma-1)T\,p},\hskip 18.49988pt(\gamma>1).$ (39)
2.3.2 Mean-reverting Gaussian premium and an induced horizon-dependent reference
We next replace the static premium by a mean-reverting Gaussian premium process, a standard reduced-form device for return predictability and intertemporal hedging (campbell2002strategic; xia2001learning). Our goal here is not to introduce additional information structure, but to obtain a closed-form, decision-time Gaussian reference for the time-averaged premium over the remaining horizon. This induces a horizon-dependent effective premium law that can be used as a controlled analytic input in numerical experiments.
OU premium dynamics and decision-time uncertainty.
Let the uncertain initial state be $\vartheta$ and set $Y_{0}=\vartheta\sim\mathcal{N}(m_{0},P_{0})$, so the decision-time law is $q=\mathcal{N}(m_{0},P_{0})$. The premium factor follows
$dY_{t}=K(\bar{y}-Y_{t})\,dt+\Xi\,dW_{t}^{Y},\hskip 18.49988ptY_{0}=\vartheta\sim\mathcal{N}(m_{0},P_{0}),$ (40)
and risky excess returns satisfy
$dR_{t}:=\frac{dS_{t}}{S_{t}}-r\mathbf{1}\,dt=BY_{t}\,dt+\Sigma^{1/2}\,dW_{t},$ (41)
allowing instantaneous correlation
$d\langle W,W^{Y}\rangle_{t}=\rho\,dt,\hskip 18.49988pt\rho\in\mathbb{R}^{d\times m}.$ (42)
Integrated premium and induced Gaussian law.
Define the integrated premium
$I_{T}:=\int_{0}^{T}Y_{s}\,ds\in\mathbb{R}^{m}.$ (43)
Since (40) is linear-Gaussian with Gaussian initial condition, $I_{T}$ is Gaussian. Its mean and covariance are
$m_{I}(T)=\mathbb{E}[I_{T}]=T\bar{y}+K^{-1}\big(I-e^{-KT}\big)\,(m_{0}-\bar{y}),$ (44)
$C_{I}(T)=\mathrm{Cov}(I_{T})$ $=K^{-1}\big(I-e^{-KT}\big)\,P_{0}\,\big(I-e^{-KT}\big)^{\top}K^{-\top}$ (45)
$\qquad+\int_{0}^{T}K^{-1}\big(I-e^{-K(T-s)}\big)\,\Xi\Xi^{\top}\,\big(I-e^{-K(T-s)}\big)^{\top}K^{-\top}\,ds.$
Notably, (44)–(45) depend only on the OU dynamics and the decision-time uncertainty $(m_{0},P_{0})$; they do not depend on the return–factor correlation $\rho$.
Horizon-averaged premium and effective Gaussian law.
Define the horizon-averaged effective premium
$\bar{\theta}_{T}:=\frac{1}{T}\,B\,I_{T}\in\mathbb{R}^{d}.$ (46)
Then
$\bar{\theta}_{T}\sim\mathcal{N}(m_{\bar{\theta}}(T),P_{\bar{\theta}}(T)),\hskip 18.49988ptm_{\bar{\theta}}(T)=\frac{1}{T}B\,m_{I}(T),\hskip 18.49988ptP_{\bar{\theta}}(T)=\frac{1}{T^{2}}B\,C_{I}(T)\,B^{\top}.$ (47)
When $\rho=0$, this induced law can be plugged directly into the static Gaussian reference of Section 2.3.1. When $\rho\neq 0$, the marginal law (47) remains valid, but constant-portfolio expected utility involves an additional cross-covariance term capturing the return–state shock linkage that generates hedging demand (campbell2002strategic; xia2001learning).
Closed-form references under constant portfolios.
Restricting to constant fractions $\pi_{t}\equiv\pi$ turns the problem into a transparent decision-time benchmark: only the integrated premium $I_{T}=\int_{0}^{T}Y_{s}\,ds$ enters the drift of $\log X_{T}^{\pi}$, while the risk term remains time-homogeneous. This yields a closed-form target that depends on the decision-time law $q=\mathcal{N}(m_{0},P_{0})$ only through the induced mean $m_{I}(T)=\mathbb{E}[I_{T}]$ (and, for CRRA, through covariances as well).
Log utility ($\gamma=1$).
With $\pi_{t}\equiv\pi$, the log-utility criterion reduces to a strictly concave quadratic in $\pi$ whose drift term depends on the OU factor only through the mean integrated premium $m_{I}(T)=\mathbb{E}\!\left[\int_{0}^{T}Y_{s}\,ds\right]$. Hence the decision-time reference depends on $q=\mathcal{N}(m_{0},P_{0})$ only through $m_{\bar{\theta}}(T)=(1/T)B\,m_{I}(T)$ (and, in particular, does not involve return–factor correlation), giving
$\pi_{q,\log}^{\mathrm{const}}(T)=\Sigma^{-1}m_{\bar{\theta}}(T)=\frac{1}{T}\,\Sigma^{-1}B\,m_{I}(T).$ (48)
CRRA ($\gamma>1$).
Define
$C_{IW}(T):=\mathrm{Cov}(I_{T},W_{T})=\int_{0}^{T}K^{-1}\big(I-e^{-K(T-s)}\big)\,\Xi\,\rho^{\top}\,ds\;\in\;\mathbb{R}^{m\times d},$ (49)
and the induced symmetric cross term
$M_{\mathrm{cross}}(T):=B\,C_{IW}(T)\,\big(\Sigma^{1/2}\big)^{\top}+\Sigma^{1/2}\,C_{IW}(T)^{\top}\,B^{\top}\;\in\;\mathbb{R}^{d\times d}.$ (50)
Then the Gaussian-$q$ decision-time reference under constant portfolios is characterized by
$\Big(\gamma T\Sigma+(\gamma-1)\big(B\,C_{I}(T)\,B^{\top}+M_{\mathrm{cross}}(T)\big)\Big)\,\pi_{q,\gamma}^{\mathrm{const}}(T)=B\,m_{I}(T),\hskip 18.49988pt(\gamma>1),$ (51)
equivalently
$\pi_{q,\gamma}^{\mathrm{const}}(T)=\Big(\gamma T\Sigma+(\gamma-1)\big(B\,C_{I}(T)\,B^{\top}+M_{\mathrm{cross}}(T)\big)\Big)^{-1}B\,m_{I}(T).$
When $\rho=0$, we have $C_{IW}(T)=0$ and $M_{\mathrm{cross}}(T)=0$, recovering the independence-case shrinkage reference.
2.4 Why dynamic programming and deep PDE surrogates break down in high-dimensional uncertain markets
This subsection explains why we do not treat classical dynamic programming (DP/HJB) or value-function-based deep PDE surrogates (PINNs / deep BSDE methods) as practical baselines in the high-dimensional uncertain markets targeted here. DP is conceptually sound in low-dimensional Markovian settings (fleming2006controlled; pham2009continuous), but two issues dominate in our regime: (i) numerically learning the value-function derivatives required for optimal policies becomes prohibitive as dimension and nonlinearity grow, and (ii) principled parameter uncertainty magnifies these difficulties.
Classical HJB: curse of dimensionality and full nonlinearity. With deterministic parameters, DP leads to an HJB for $V(t,x,y)$ (fleming2006controlled). Grid-based solvers scale exponentially in the state dimension (bellman1961adaptive; kushner2001numerical). In portfolio problems with $d$ assets and $m$ factors, the natural state already has dimension $m+2$, so even modest discretizations require $N^{m+2}$ grid points. Moreover, constraints, transaction costs, and non-affine dynamics typically yield fully nonlinear HJBs, where stable monotone schemes are delicate even in moderate dimension and become impractical in the regime we target (kushner2001numerical).
Deep PDE surrogates: fewer grids, same derivative bottleneck. PINNs and deep BSDE methods replace grids with neural approximators trained on sampled points/paths (raissi2019physics; sirignano2018dgm; han2018solving; beck2019machine), but for fully nonlinear portfolio HJBs they remain value-function-based: they must implicitly learn high-dimensional gradients/Hessians and, crucially, mixed sensitivities (e.g. $V_{xy}$) that drive intertemporal hedging. In practice this induces nonconvex, ill-conditioned objectives (due to control suprema and nonlinear derivative dependence) and training signals that do not reliably control the specific derivative components needed for stable hedging demands in high dimension.
Latent parameter uncertainty: belief-state blowup and $\theta$-blind aggregation. A principled DP treatment augments the state with a posterior/belief over parameters, leading to a value function $V(t,x,y,\Pi)$ on a space of measures in general (bensoussan1985optimal; pham2017dynamic). Even when finite-dimensional conjugate summaries exist, the enlarged HJB is substantially harder. For deep surrogates, uncertainty either requires solving many $\theta$-conditional problems (expensive) or treating $\theta$ as an extra input (higher effective dimension, worse conditioning). In our deployable $\theta$-blind setting, a single policy must perform well under $\theta\sim q$, coupling heterogeneous models and potentially causing high-variance gradients and cancellation across parameter draws.
We therefore avoid value-function PDE/BSDE baselines in this regime and instead work with a $q$-aggregated Pontryagin stationarity condition and projection map, estimating expectations over $\theta\sim q$ via Monte Carlo inside the simulator. While $\theta$-conditional PMP objects can still be computed under frozen-$\theta$ simulations for inspection, our deployable target and guarantees are expressed in terms of $q$-aggregated stationarity, motivating the simulation-based methods in Section 3.
3 Pontryagin–Guided Policy Optimization under Latent Parameter Uncertainty
We study the fixed-$q$ ex–ante portfolio choice problem of Section 2 under latent parameter uncertainty $\theta\sim q$. The investor must deploy a $\theta$-blind policy (Remark 1), so the control can depend on observable states $(t,X_{t},Y_{t})$ but cannot take $\theta$ as an input. We restrict attention to Markov feedback policies parameterized by a neural network $\pi_{\varphi}(t,x,y)$.
Our solution approach follows a two-stage pipeline:
  • Stage 1 (PG–DPO). We perform stochastic gradient ascent on the ex–ante objective $J(\varphi)=\mathbb{E}[U(X_{T}^{\pi_{\varphi},\theta})]$, sampling $\theta$ only inside the simulator while keeping $\pi_{\varphi}$ deployable and $\theta$-blind.
  • Stage 2 (Pontryagin projection). Under a warm-up policy, we estimate Pontryagin sensitivity objects by BPTT (conditionally on frozen $\theta$), aggregate them across $\theta\sim q$, and construct a single deployable portfolio by projecting onto the aggregated first-order condition (20).
A practical subtlety is that the $q$-aggregated Pontryagin condition involves mixed moments across $\theta$ (products of $\theta$-dependent costates and $\theta$-dependent coefficients). In moderate to high dimensions, these quantities can be statistically noisy under finite Monte Carlo budgets. In our implementation, the main stabilization mechanisms are (i) estimating stage 2 objects under a warm-up policy (two-time-scale stabilization), (ii) computing the same projection in a residual/control-variate form (Section 3.3.1), and (iii) amortizing projection via interactive distillation (Section 3.3.2).
Section 3.1 reviews baseline PG–DPO and the conditional BPTT–PMP correspondence. Section 3.2 develops the stage 2 $q$-aggregated projection under latent $\theta$, together with a residual-based policy-gap guarantee. Section 3.3 records two practical couplings between stage 1 and stage 2 (residual form and interactive distillation).
3.1 PG–DPO as stochastic gradient ascent and BPTT–PMP correspondence
Setup and objectives (frozen $\theta$, deployable $\theta$-blind feedback).
A latent parameter $\theta\in\Theta$ is sampled from a fixed law $q(d\theta)$ inside the simulator and kept frozen along each simulated trajectory. A deployable portfolio policy is a $\theta$-blind Markov feedback rule represented by a neural network
$\pi_{\varphi}:\ [0,T]\times(0,\infty)\times\mathbb{R}^{m}\to\mathbb{R}^{d},\hskip 18.49988pt(t,x,y)\mapsto\pi_{\varphi}(t,x,y),\hskip 18.49988pt\varphi\in\mathbb{R}^{p},$ (52)
which does not take $\theta$ as an input. For a fixed frozen $\theta$, the $\theta$-conditional state is $(X_{t}^{\pi,\theta},Y_{t}^{\theta})_{t\in[0,T]}$ and evolves as
$\frac{dX_{t}^{\pi,\theta}}{X_{t}^{\pi,\theta}}$ $=\Big(r+\pi_{t}^{\top}b\big(Y_{t}^{\theta},\theta\big)\Big)\,dt+\pi_{t}^{\top}\sigma\big(Y_{t}^{\theta},\theta\big)\,dW_{t},\hskip 18.49988ptX_{0}=x>0,$ (53)
$dY_{t}^{\theta}$ $=a\big(Y_{t}^{\theta},\theta\big)\,dt+\beta\big(Y_{t}^{\theta},\theta\big)\,dW_{t}^{Y},\hskip 18.49988ptY_{0}=y\in\mathbb{R}^{m}.$ (54)
For each fixed $\theta$ we evaluate $\pi_{\varphi}$ by the conditional objective
$J^{\theta}(\varphi):=\mathbb{E}\big[U\big(X_{T}^{\pi_{\varphi},\theta}\big)\,\big|\,\theta\big],$ (55)
where the expectation is over Brownian paths in (53)–(54). The fixed-$q$ ex–ante objective is
$J(\varphi):=\mathbb{E}_{\theta\sim q}\big[J^{\theta}(\varphi)\big]=\mathbb{E}\Big[U\big(X_{T}^{\pi_{\varphi},\theta}\big)\Big],$ (56)
where the last expectation is joint over $\theta\sim q$ and $(W,W^{Y})$. Thus $\sup_{\varphi}J(\varphi)$ is a standard stochastic optimization problem: $\theta$ is sampled inside the simulator while the policy remains $\theta$-blind.
Discretization, sampling over $\theta$, and baseline PG–DPO update.
We discretize $[0,T]$ into $N$ steps of length $\Delta t$ and approximate (53)–(54) by an Euler scheme. For episode $i$ we denote the discrete state by $(X_{k}^{(i)},Y_{k}^{(i)})_{k=0,\dots,N}$ and write $\theta^{(i)}$ for the frozen parameter used to generate that simulated environment. Given $\pi_{\varphi}$ and Brownian increments, the mapping
$\big(x^{(i)},y^{(i)},\theta^{(i)},\{\Delta W_{k}^{(i)},\Delta W_{k}^{Y,(i)}\}_{k=0}^{N-1},\varphi\big)\longmapsto U\big(X_{N}^{(i)}\big)$
is a finite computational graph, so automatic differentiation computes exact discrete gradients $\nabla_{\varphi}U(X_{N}^{(i)})$.
A typical PG–DPO update samples a mini-batch of initial states $\{(t_{0}^{(i)},x_{0}^{(i)},y_{0}^{(i)})\}_{i=1}^{M}$ from a user-chosen training distribution $\nu$ on $0,T)\times(0,\infty)\times\mathbb{R}^{m}$, samples $\theta\sim q$ inside the simulator (unseen by the policy) and holds it frozen for the update, and simulates forward:
$Y_{k+1}^{(i)}$ $=Y_{k}^{(i)}+a\big(Y_{k}^{(i)},\theta\big)\Delta t^{(i)}+\beta\big(Y_{k}^{(i)},\theta\big)\Delta W_{k}^{Y,(i)},$
$X_{k+1}^{(i)}$ $=X_{k}^{(i)}+X_{k}^{(i)}\Big(r+\pi_{\varphi}\big(t_{k}^{(i)},X_{k}^{(i)},Y_{k}^{(i)}\big)^{\top}b\big(Y_{k}^{(i)},\theta\big)\Big)\Delta t^{(i)}$
$\hphantom{={}}{}+X_{k}^{(i)}\,\pi_{\varphi}\big(t_{k}^{(i)},X_{k}^{(i)},Y_{k}^{(i)}\big)^{\top}\sigma\big(Y_{k}^{(i)},\theta\big)\Delta W_{k}^{(i)},$
starting from $X_{0}^{(i)}=x_{0}^{(i)}$, $Y_{0}^{(i)}=y_{0}^{(i)}$. The episode reward is
$J^{(i)}(\varphi):=U\big(X_{N}^{(i)}\big),$ (57)
and BPTT computes $\nabla_{\varphi}J^{(i)}(\varphi)$. The policy parameters are then updated (e.g. by Adam) as
$\varphi\leftarrow\varphi+\alpha\,\frac{1}{M}\sum_{i=1}^{M}\nabla_{\varphi}J^{(i)}(\varphi).$ (58)
Sampling $\theta$ independently per episode (i.e. $\theta^{(i)}\sim q$) or sampling one $\theta\sim q$ per update and reusing it across the batch both yield unbiased stochastic gradients for $J(\varphi)$.
Pathwise costates from BPTT and the (conditional) BPTT–PMP correspondence.
BPTT returns not only $\nabla_{\varphi}J^{(i)}(\varphi)$ but also pathwise sensitivities with respect to intermediate state variables, which coincide with discrete-time adjoint variables (costates) in the sense of Pontryagin. For a single episode (suppressing $i$ and $\theta$ in notation), define the pathwise wealth costate
$p_{k}:=\frac{\partial U(X_{N})}{\partial X_{k}},\hskip 18.49988ptk=0,\dots,N,$ (59)
and the additional pathwise sensitivity objects used in projected controls:
$p_{x,k}:=\frac{\partial p_{k}}{\partial X_{k}},\hskip 18.49988ptp_{y,k}:=\frac{\partial p_{k}}{\partial Y_{k}},\hskip 18.49988ptk=0,\dots,N.$ (60)
Theorem 2 (BPTT–PMP correspondence (conditional on $\theta$, uniform on compacts)).
Fix $\theta\in\Theta$ and assume standard regularity conditions ensuring (i) well-posedness of the $\theta$-conditional forward SDE (53)–(54) under the $\theta$-blind policy $\pi_{\varphi}$ and (ii) well-posedness of the associated $\theta$-conditional stochastic maximum principle (adjoint) system. Let $(p_{t}^{\theta},p_{x,t}^{\theta},p_{y,t}^{\theta})$ denote the resulting continuous-time Pontryagin objects under $\pi_{\varphi}$ (and, in smooth Markov regimes, the corresponding spatial derivatives of the decoupling field). Let $(p_{k},p_{x,k},p_{y,k})$ be the discrete pathwise quantities computed by BPTT for the Euler discretization with step $\Delta t$, as defined in (59)–(60).
Then, as $\Delta t\to 0$, the BPTT-induced discrete adjoints converge to their continuous-time counterparts in an appropriate mean-square sense (along trajectories). Moreover, for any compact set $K\subset\Theta$, the constants in the convergence bounds can be chosen uniformly for all $\theta\in K$.
Proof.
See Appendix B. ∎
Across $\theta\sim q$, these Pontryagin objects form a $\theta$-indexed family. Baseline PG–DPO trains against the ex–ante objective (56) by repeatedly sampling $\theta\sim q$ inside the simulator, while the deployable policy remains $\theta$-blind.
3.2 Projected PG–DPO under latent $\theta$: $q$-aggregated projection and a residual-based policy-gap bound
Stage 2 is a projection step: given a warm-up deployable $\theta$-blind feedback policy $\pi^{\mathrm{warm}}=\pi_{\varphi^{\mathrm{warm}}}$ (from stage 1), we estimate $\theta$-conditional Pontryagin sensitivity objects by BPTT/Monte Carlo under frozen $\theta\sim q$, aggregate them across $\theta$, and construct a deployable $\theta$-blind policy by projecting onto the $q$-aggregated Pontryagin stationarity condition derived in Section 2.2. The main point is that the aggregated first-order condition is affine in the portfolio control; hence it induces a statewise linear system and, on a suitable working domain, a concrete projection map from estimated Pontryagin objects to a portfolio rule.
Working domain and norms. Fix a measurable working state domain $D\subset[0,T]\times(0,\infty)\times\mathbb{R}^{m}$ (e.g. a training/evaluation band) and a reference measure $\mu$ on $D$ (e.g. an empirical state distribution induced by rollouts). For $h:D\to\mathbb{R}^{n}$ we write
$\|h\|_{L^{2}(\mu)}:=\Big(\int_{D}\|h(z)\|^{2}\,\mu(dz)\Big)^{1/2},\hskip 18.49988ptz=(t,x,y),$
and for $\theta$-indexed families (used when tracking frozen-$\theta$ quantities in analysis/inspection),
$\|f\|_{L^{2}(q\otimes\mu)}:=\bigg(\int_{\Theta}\int_{D}\|f^{\theta}(z)\|^{2}\,\mu(dz)\,q(d\theta)\bigg)^{1/2}.$ (61)
Mixed-moment $q$-aggregation under a warm-up policy.
By Theorem 1, any locally optimal interior deployable $\theta$-blind policy $\pi^{\star,\mathrm{blind}}$ for the fixed-$q$ ex–ante problem satisfies the $q$-aggregated stationarity condition (20). In the portfolio Hamiltonian (15), this stationarity is equivalent to a statewise linear system and hence to the projected form (25) on the working domain (under invertibility of the aggregated curvature term). P–PGDPO constructs a practical approximation of this projection by estimating the relevant aggregated Pontryagin objects under a fixed warm-up policy $\pi^{\mathrm{warm}}=\pi_{\varphi^{\mathrm{warm}}}$.
Fix a query state $z=(t,x,y)\in D$ and a frozen parameter $\theta$. We simulate trajectories under $\pi^{\mathrm{warm}}$ and compute pathwise Pontryagin sensitivity objects by autodiff/BPTT; averaging over $M_{\mathrm{MC}}$ trajectories yields Monte Carlo estimates
$\widehat{p}_{t}^{\theta}(z),\hskip 18.49988pt\widehat{p}_{x,t}^{\theta}(z),\hskip 18.49988pt\widehat{p}_{y,t}^{\theta}(z).$ (62)
Using these, define the $\theta$-conditional estimated projection inputs
$\widehat{A}_{t}^{\theta}(t,x,y)$ $:=x\,\widehat{p}_{x,t}^{\theta}(t,x,y)\,\Sigma(y,\theta)\in\mathbb{R}^{d\times d},$ (63)
$\widehat{G}_{t}^{\theta}(t,x,y)$ $:=\widehat{p}_{t}^{\theta}(t,x,y)\,b(y,\theta)+\Sigma_{SY}(y,\theta)\,\widehat{p}_{y,t}^{\theta}(t,x,y)\in\mathbb{R}^{d}.$ (64)
Aggregating across $\theta\sim q$ (approximated in practice by sampling $M_{\theta}$ frozen parameters) gives
$\widehat{A}_{t}(t,x,y)$ $:=\mathbb{E}_{\theta\sim q}\Big[\widehat{A}_{t}^{\theta}(t,x,y)\Big],$ (65)
$\widehat{G}_{t}^{\mathrm{mix}}(t,x,y)$ $:=\mathbb{E}_{\theta\sim q}\Big[\widehat{G}_{t}^{\theta}(t,x,y)\Big].$ (66)
Whenever $\widehat{A}_{t}(t,x,y)$ is invertible on $D$, we obtain the mixed-moment projected policy
$\widehat{\pi}^{\mathrm{agg,mix}}(t,x,y):=-\,\widehat{A}_{t}(t,x,y)^{-1}\,\widehat{G}_{t}^{\mathrm{mix}}(t,x,y).$ (67)
Residual diagnostic and a slab-wise small-gain policy-gap bound.
To connect the projected policy (67) to a locally optimal deployable $\theta$-blind policy, we measure how well the warm-up policy satisfies the population mixed-moment aggregated stationarity. Let $(A_{\pi},G_{\pi}^{\mathrm{mix}})$ denote the mixed-moment $q$-aggregated projection inputs induced by a policy $\pi$ (i.e. the objects in (23) evaluated using the $\theta$-conditional Pontryagin objects generated by $\pi$). Define the warm-up aggregated stationarity residual on $D$ by
$r^{\mathrm{warm}}_{\mathrm{FOC,mix}}(t,x,y):=A_{\pi^{\mathrm{warm}}}(t,x,y)\,\pi^{\mathrm{warm}}(t,x,y)+G_{\pi^{\mathrm{warm}}}^{\mathrm{mix}}(t,x,y),\hskip 18.49988pt\varepsilon_{\mathrm{warm}}^{\mathrm{mix}}:=\big\|r^{\mathrm{warm}}_{\mathrm{FOC,mix}}\big\|_{L^{2}(\mu)}.$ (68)
In practice we monitor the estimator $\widehat{r}^{\mathrm{warm}}_{\mathrm{FOC,mix}}:=\widehat{A}_{t}\,\pi^{\mathrm{warm}}+\widehat{G}_{t}^{\mathrm{mix}}$ computed from the same BPTT/Monte Carlo pipeline.
A technical point is that a global small-gain condition of the form $C_{1}<1$ can be overly restrictive. Following the slab-wise philosophy in our prior PGDPO analysis (e.g. huh2025breaking), we default to a time-slab decomposition of the working domain and close the warm-up gap on each short slab. Concretely, assume $D$ carries a time coordinate and fix a partition $0=t_{0}<t_{1}<\cdots<t_{K}=T$ with slab lengths $\tau_{k}:=t_{k}-t_{k-1}$. Let
$D_{k}:=D\cap([t_{k-1},t_{k}]\times\mathcal{S}),\hskip 18.49988pt\mu_{k}:=\mu|_{D_{k}},\hskip 18.49988pt\|f\|_{k}:=\|f\|_{L^{2}(\mu_{k})}.$
We write $T(\pi):=-A_{\pi}^{-1}G_{\pi}^{\mathrm{mix}}$ for the (population) $q$-aggregated projection map. Theorem 3 below shows that, under a mild slab-wise local stability regime (i.e. a short-time contraction of $T$ on each $D_{k}$), small residual implies that the projected policy is close (in $L^{2}(\mu)$) to a locally optimal deployable $\theta$-blind policy, up to discretization/Monte Carlo error. The proof combines a projection-map stability bound (Appendix C.1) with a slab-wise closure (Appendix C.2), in the same spirit as the slab analyses used in huh2025breaking.
Theorem 3 (Residual-based ex–ante $\theta$-blind policy-gap bound for P–PGDPO (mixed-moment, deployable, slab-wise local)).
Assume the uniform invertibility/stability conditions of Proposition 1 (Appendix C.1) hold on $D$ for the relevant aggregated curvature terms and for the estimator perturbations constructed under $\pi^{\mathrm{warm}}$.
Let $\pi^{\star,\mathrm{blind}}$ be a locally optimal interior deployable $\theta$-blind policy for the fixed-$q$ ex–ante problem. Assume there exists a neighborhood $\mathcal{U}$ of $\pi^{\star,\mathrm{blind}}$ in $L^{2}(\mu)$ such that for all $\pi\in\mathcal{U}$,
$\|A_{\pi}^{-1}\|_{L^{\infty}(D)}\leq\kappa,\hskip 18.49988pt\|G_{\pi}^{\mathrm{mix}}\|_{L^{\infty}(D)}\leq M_{G},$
and assume the slab-wise Lipschitz gain of Appendix C.2 holds: there exist constants $\bar{L}_{A},\bar{L}_{G}>0$ such that for every slab $D_{k}$ and all $\pi_{1},\pi_{2}\in\mathcal{U}$,
$\|A_{\pi_{1}}-A_{\pi_{2}}\|_{k}\leq\bar{L}_{A}\,\tau_{k}^{1/2}\,\|\pi_{1}-\pi_{2}\|_{k},\hskip 18.49988pt\|G_{\pi_{1}}^{\mathrm{mix}}-G_{\pi_{2}}^{\mathrm{mix}}\|_{k}\leq\bar{L}_{G}\,\tau_{k}^{1/2}\,\|\pi_{1}-\pi_{2}\|_{k}.$ (69)
Define
$\rho(\tau):=\big(\kappa\bar{L}_{G}+\kappa^{2}M_{G}\bar{L}_{A}\big)\tau^{1/2},\hskip 18.49988pt\rho_{*}:=\max_{1\leq k\leq K}\rho(\tau_{k}).$ (70)
Assume the slab partition is chosen so that $\rho_{*}<1$.
Let $\widehat{\pi}^{\mathrm{agg,mix}}$ be the mixed-moment projected policy (67) computed from BPTT/Monte Carlo estimates under $\pi^{\mathrm{warm}}$, and let $\varepsilon_{\mathrm{warm}}^{\mathrm{mix}}$ be the population residual (68). Then there exists $C_{2}>0$ such that
$\big\|\widehat{\pi}^{\mathrm{agg,mix}}-\pi^{\star,\mathrm{blind}}\big\|_{L^{2}(\mu)}\;\leq\;\frac{\rho_{*}\kappa}{1-\rho_{*}}\,\varepsilon_{\mathrm{warm}}^{\mathrm{mix}}\;+\;C_{2}\,\delta_{\mathrm{BPTT}}(\Delta t,M_{\mathrm{MC}},M_{\theta}).$ (71)
Moreover, under the perturbative regime of Proposition 1, one may take for example $C_{2}:=2\kappa+4\kappa^{2}M_{G}$.
Proof.
See Appendix C.3. ∎
3.3 Coupling stage 1 and stage 2: residual projection and interactive distillation
We keep the ex–ante objective (56) and the $\theta$-blind deployability constraint throughout. Stage 2 is not a separate optimization problem: it reuses the current stage 1 policy as a warm-up control under which the (costate-based) projection ingredients are estimated, and then applies a $q$-aggregated Pontryagin projection as a post-processing map.
This subsection records two couplings between the two stages, each with a distinct role. First, we implement the projected rule in a residual (control-variate) form, which is algebraically equivalent to the direct projection but typically reduces Monte Carlo variance and improves numerical stability in high dimensions. Second, we use the projected output as a teacher signal via interactive distillation. Beyond acting as an optimization aid, distillation serves an amortization purpose: stage 2 projection can be accurate but Monte-Carlo intensive, whereas a distilled student policy can approximate the projected rule with a single forward pass at stage 1 inference cost.
3.3.1 Control-variate (residual) form of the projected rule
Recall the mixed-moment projected rule (67). In high dimensions, Monte Carlo noise in the projection inputs can be non-negligible, and solving a linear system with $\widehat{A}_{t}$ can amplify this noise. A convenient stabilization is to compute the same projected rule in a residual form around the warm-up policy $\pi_{\varphi^{\mathrm{warm}}}$.
Define the $\theta$-conditional residual (under frozen-$\theta$ simulations)
$\widehat{r}_{\mathrm{FOC}}^{\theta}(t,x,y):=\widehat{A}_{t}^{\theta}(t,x,y)\,\pi_{\varphi^{\mathrm{warm}}}(t,x,y)+\widehat{G}_{t}^{\theta}(t,x,y),$ (72)
and the aggregated residual (the quantity we actually solve against)
$\widehat{r}_{\mathrm{FOC}}(t,x,y):=\widehat{A}_{t}(t,x,y)\,\pi_{\varphi^{\mathrm{warm}}}(t,x,y)+\widehat{G}_{t}^{\mathrm{mix}}(t,x,y).$ (73)
Whenever $\widehat{A}_{t}(t,x,y)$ is invertible, the projected rule admits the identity
$\widehat{\pi}^{\mathrm{agg,mix}}(t,x,y)=\pi_{\varphi^{\mathrm{warm}}}(t,x,y)-\widehat{A}_{t}(t,x,y)^{-1}\,\widehat{r}_{\mathrm{FOC}}(t,x,y),$ (74)
which is an algebraic rewriting of (67) (hence it does not change the target). Its practical value is variance reduction: when the warm-up policy is already close to a projected fixed point on the working domain, the residual $\widehat{r}_{\mathrm{FOC}}$ tends to be small, and it often concentrates faster because the ingredients entering $\widehat{A}_{t}\pi_{\varphi^{\mathrm{warm}}}$ and $\widehat{G}_{t}^{\mathrm{mix}}$ are computed from the same Monte Carlo pool and partially cancel.
3.3.2 Interactive distillation: projection-guided training and amortized deployment
Let $\pi_{\varphi}$ be the trainable stage 1 policy network. At intermittent refresh times, we freeze a lagged copy $\pi_{\varphi^{-}}$ and run stage 2 under $\pi_{\varphi^{-}}$ to construct a $q$-aggregated projected teacher. This coupling serves two purposes. During training it provides projection-guided targets that can stabilize and accelerate stage 1 optimization; after training it amortizes the expensive projection by distilling it into a fast deployable policy network.
In residual form (74), the teacher is the $\theta$-blind map
$\widehat{\pi}^{\mathrm{agg,mix}}_{\varphi^{-}}(t,x,y):=\pi_{\varphi^{-}}(t,x,y)-\big(\widehat{A}_{t}^{\varphi^{-}}(t,x,y)\big)^{-1}\,\widehat{r}_{\mathrm{FOC}}^{\varphi^{-}}(t,x,y),$ (75)
where $\widehat{r}_{\mathrm{FOC}}^{\varphi^{-}}:=\widehat{A}_{t}^{\varphi^{-}}\pi_{\varphi^{-}}+\widehat{G}_{t}^{\mathrm{mix},\varphi^{-}}$ is computed using the mixed-moment $q$-aggregation under the lagged policy. We then train $\pi_{\varphi}$ by combining the original ex–ante objective with a proximity term to this teacher on the working domain:
$\max_{\varphi}\;J(\varphi)\;-\;\lambda\,\mathbb{E}_{(t,x,y)\sim\mu}\Big[\big\|\pi_{\varphi}(t,x,y)-\mathrm{stopgrad}\big(\widehat{\pi}^{\mathrm{agg,mix}}_{\varphi^{-}}(t,x,y)\big)\big\|^{2}\Big],$ (76)
where $\mu$ is the working-domain sampling measure and $\lambda\geq 0$ controls the strength of projection guidance. The operator $\mathrm{stopgrad}(\cdot)$ indicates that gradients are not propagated through stage 2; once computed from $\pi_{\varphi^{-}}$, the teacher is treated as fixed.
In practice, $\varphi^{-}$ and $\widehat{\pi}^{\mathrm{agg,mix}}_{\varphi^{-}}$ are refreshed on a slower timescale than the stage 1 gradient steps: we hold $\varphi^{-}$ fixed for several updates of $\varphi$ under (76), then set $\varphi^{-}\leftarrow\varphi$ and recompute the teacher. A practical schedule is to start with $\lambda=0$ (pure PG–DPO) and increase $\lambda$ only after basic projection checks on the working domain (e.g., residual magnitudes and curvature/denominator stability) indicate that the stage 2 map has become reliable. Moreover, to avoid injecting noisy teacher targets early in training or on pathological regions of the domain, we may apply projection guidance only on states where the projection checks certify reliability (an “adaptive teacher selection”); implementation details are deferred to the appendix (Appendix D).
4 Breaking the Dimensional Barrier under Drift Uncertainty
This section instantiates the decision-time static Gaussian drift-uncertainty benchmark in Section 2.3.1 and uses its closed-form constant-portfolio $q$-reference as an analytic target. Nature draws a fixed latent drift $\theta\sim q$ at $t=0$ and keeps it constant over $[0,T]$, while the investor cannot observe $\theta$ and must deploy a single $\theta$-blind policy under an ex–ante CRRA objective. Because the benchmark admits a transparent decision-time reference, we can measure accuracy directly via decision-time RMSE, rather than relying only on realized utility.
Our goal is to test whether Pontryagin-guided learning and projection remain stable as the number of assets grows. We generate APT-style covariance structures and sweep dimensions $d\in\{5,10,50,100\}$ under both aligned uncertainty ($P=s\Sigma$) and a misaligned geometry that rotates uncertainty away from market risk directions. We compare Stage 1 (PG–DPO) to Stage 2 (Pontryagin projection) (and, when applicable, amortized variants via interactive distillation) under matched simulation budgets that scale linearly with $d$.
4.1 Benchmark market and evaluation protocol
This subsection fixes the benchmark and evaluation protocol used in Section 4. Our goal is to provide controlled evidence that the proposed two-stage pipeline remains computationally stable and accurate as the number of assets $d$ grows under decision-time parameter uncertainty. The aligned vs. misaligned uncertainty geometries serve as two representative stress-test regimes; the main message is scalability under uncertainty rather than any specific choice of $P$.
$\theta$-blind deployability (and what uses $\theta$). Throughout Section 4, all reported policies are deployable and $\theta$-blind: the control is a function of observable state only (here, decision-time evaluation uses $t=0$ and $X_{0}$), and never takes the realized latent premium $\theta$ as an input. The latent $\theta\sim q$ is sampled only inside the simulator to generate trajectories and to form Monte Carlo averages that approximate $q$-expectations (notably in Stage 2 projection). Any $\theta$-indexed objects (when referenced elsewhere) are used only for offline diagnostics and are not part of the deployable decision rule.
Static decision-time uncertainty benchmark. We adopt the static Gaussian drift-uncertainty market of Section 2.3.1, i.e., (26) with (35). Equivalently, we simulate
$\frac{dS_{t}}{S_{t}}=r\,\mathbf{1}\,dt+\theta\,dt+\Sigma^{1/2}dW_{t},\hskip 18.49988pt\theta\sim\mathcal{N}(m,P),$
where the latent premium $\theta$ is drawn once at time 0 and kept fixed over $[0,T]$. The deployable policy is $\theta$-blind and interacts with $q$ only through sampling $\theta$ inside the simulator.
APT-style factor construction of $(m,\Sigma)$. We construct the mean premium and covariance via a low-dimensional factor representation. Let $W^{f}$ be a $k_{\Sigma}$-dimensional Brownian motion (factor shocks) and $W^{\varepsilon}$ a $d$-dimensional Brownian motion (idiosyncratic shocks), independent of $W^{f}$. We write excess returns as
$dR_{t}:=\frac{dS_{t}}{S_{t}}-r\mathbf{1}\,dt=\theta\,dt+B\,\Sigma_{f}^{1/2}\,dW_{t}^{f}+\mathrm{diag}(\sqrt{D})\,dW_{t}^{\varepsilon},$ (77)
with $B\in\mathbb{R}^{d\times k_{\Sigma}}$, $\Sigma_{f}\succ 0$, and $D\in(0,\infty)^{d}$. This implies
$\Sigma=B\,\Sigma_{f}\,B^{\top}+\mathrm{diag}(D)\;=\;FF^{\top}+\mathrm{diag}(D),$ (78)
where $F:=B\,\mathrm{chol}(\Sigma_{f})$. We generate the mean premium in an APT-like form by drawing a factor price vector $\lambda_{m}\in\mathbb{R}^{k_{\Sigma}}$ and setting
$m:=B\,\lambda_{m}.$ (79)
One-shot generation and fairness across methods. For each dimension $d$, we generate a single market instance $(B,\Sigma_{f},D,\lambda_{m})$ once (using a fixed random seed) and hold it fixed across all algorithmic comparisons and MC-budget variants. Within a fixed $d$, we change only the uncertainty covariance $P$ (aligned vs. misaligned and the scale $s$ below). This isolates algorithmic effects from instance-to-instance randomness and makes the scaling comparisons controlled.
Uncertainty regimes (aligned vs. misaligned). We consider two geometries for the drift-uncertainty covariance $P$, controlled by a scalar magnitude $s>0$.
Aligned: uncertainty shares market risk directions,
$P=s\,\Sigma,\hskip 18.49988pts>0.$ (80)
Misaligned: uncertainty factors are rotated away from the market factor space,
$P=\widetilde{B}\,\widetilde{\Sigma}_{f}\,\widetilde{B}^{\top}+s\,\mathrm{diag}(D),$ (81)
where $\widetilde{B}$ is generated independently of $B$ (or explicitly orthogonalized against the span of $B$ to enforce large principal angles). The factor term is rescaled so that its overall magnitude matches the aligned case under the same $s$ (e.g., by matching $\mathrm{tr}(P)$ or $\|P\|_{F}$ up to the shared diagonal component). This geometry increases heterogeneity across $\theta\sim q$ and makes mixed-moment estimation and subsequent linear-algebra steps more fragile, providing a stringent scalability test.
Experiment grid and simulation budgets. We vary the number of assets over $d\in\{5,10,50,100\}$ and sweep three uncertainty magnitudes $s\in\{10^{-3},10^{-2},10^{-1}\}$, for both aligned and misaligned geometries. To keep Monte Carlo noise comparable across dimensions, we use linear-in-$d$ sampling budgets: a base regime with $N_{\mathrm{MC}}=100\cdot d$ paths and a high regime with $N_{\mathrm{MC}}=400\cdot d$ (where $N_{\mathrm{MC}}$ denotes the per-update or per-estimator path budget, depending on the stage). All methods share the same discretization scheme (Euler) and action constraints; implementation details (network architecture, optimizer settings, and exact sampling conventions for Stage 1 vs. Stage 2) are reported in the implementation appendix and code release.
Analytic reference and decision-time evaluation. In the static Gaussian benchmark, the analytic decision-time reference under constant portfolios is available in closed form. We use this closed-form rule only as an external decision-time target for evaluation; training does not impose the constant-portfolio restriction, and all methods learn from simulated trajectories over $[0,T]$ under the same $\theta$-blind constraint. For $\gamma>1$ we use the CRRA reference (38) and, for $\gamma=1$, the log-utility reference (31). We evaluate each method at $t=0$ on a fixed grid $\{(X_{0}^{(i)},T^{(i)})\}_{i=1}^{N_{\mathrm{eval}}}$ and report RMSE to the analytic reference:
$\mathrm{RMSE}(u_{0},\pi_{q,\gamma}^{\mathrm{const}}):=\bigg(\frac{1}{N_{\mathrm{eval}}}\sum_{i=1}^{N_{\mathrm{eval}}}\big\|u_{0}(X_{0}^{(i)},T^{(i)})-\pi_{q,\gamma}^{\mathrm{const}}(T^{(i)})\big\|^{2}\bigg)^{1/2},$ (82)
where $u_{0}(\cdot)$ denotes the decision-time action prescribed by the method (deployable $\theta$-blind output). With the benchmark fixed and with $(m,\Sigma,P)$ constructed as in (79)–(81), the remaining subsections compare baseline Stage 1 PG–DPO, post-hoc Stage 2 P–PGDPO projection, and interactive distillation under matched simulation budgets.
4.2 High-dimensional CRRA benchmark: projection and amortization
Mixed-moment estimation and a decoupling approximation. A practical issue throughout our experiments (both aligned and misaligned) is the estimation of mixed moments across the latent parameter, such as $\mathbb{E}_{\theta\sim q}[p_{t}^{\theta}(z)\,\theta]$ (and analogous products entering $\widehat{G}_{t}^{\mathrm{mix}}$), because the costate $p_{t}^{\theta}(z)$ is $\theta$-dependent and high-dimensional, and finite-sample covariance between $p_{t}^{\theta}$ and $\theta$ can lead to large Monte-Carlo variance once the subsequent linear solve is applied. For numerical stability and a uniform protocol across geometries, we therefore use a simple decoupling (independence) approximation for these mixed moments,
$\mathbb{E}_{\theta\sim q}[p_{t}^{\theta}(z)\,\theta]\;\approx\;\mathbb{E}_{\theta\sim q}[p_{t}^{\theta}(z)]\,\mathbb{E}_{\theta\sim q}[\theta],$
(and similarly for other mixed products), which is exact when the relevant Pontryagin objects are effectively $\theta$-invariant and is accurate whenever $\mathrm{Cov}_{q}(p_{t}^{\theta}(z),\theta)$ is small relative to marginal scales. While this approximation is most valuable under misalignment—where direction mixing can amplify mixed-moment noise—it also performs well in aligned regimes (where mixed moments are typically easier to estimate), and in the CRRA benchmark below it does not alter the qualitative scaling conclusions: projection remains stable, and the observed misaligned degradation is consistent with residual growth and curvature mismatch rather than catastrophic mixed-moment blow-ups.  [^1] We note, however, that in extreme uncertainty/misalignment—where $\theta$–costate dependence becomes pronounced—the decoupling can break down, in which case one should revert to full mixed-moment estimation (possibly with larger budgets and/or regularized/certified projection).
Figure 1: Decision-time RMSE at $t=0$ versus dimension $d$ (log scale), summarized by a tail median over the late-training window (computed from the last evaluation snapshots). Rows: uncertainty magnitude $s\in\{10^{-3},10^{-2},10^{-1}\}$. Columns: aligned vs. misaligned geometry. Curves compare Stage 1 (deployable) and Stage 2 (post-hoc projection), with and without interactive distillation. Solid vs. dashed lines correspond to MC base ($100\cdot d$) vs. high ($400\cdot d$) budgets.
Protocol and summary statistic. We consider the CRRA benchmark with $\gamma=2$ under Gaussian drift uncertainty $q$ and evaluate against the analytic constant $q$-reference (38). We track (i) the Monte-Carlo objective estimate $\widehat{J}$ during training and (ii) the decision-time error at $t=0$ via RMSE (82). Because stochastic optimization produces non-monotone and noisy RMSE curves, we summarize each condition by a robust tail median: the median RMSE over the final evaluation snapshots in the late-training window. Unless stated otherwise, the projection/teacher direction uses the mixed-moment ($p_{\theta}$) aggregation.
What is compared in Figure 1. Stage 1 (PG–DPO; Section 3.1) trains a deployable $\theta$-blind policy $\pi_{\varphi}$ by maximizing $\widehat{J}$ via pathwise gradients. Stage 2 (P–PGDPO; Section 3.2) applies a $q$-aggregated Pontryagin projection to a Stage 1 checkpoint; we use the residual form of Section 3.3.1. Interactive distillation (Section 3.3.2) treats the Stage 2 projected control as a teacher signal and amortizes it back into a deployable Stage 1 policy network.
Thus Figure 1 separates projection quality (Stage 2: post-hoc projected, still $\theta$-blind) from amortized deployable quality (Stage 1 distilled: single forward pass).
Stage 2 projection versus amortization: scaling with dimension. Aligned geometry. For small and moderate uncertainty ($s=10^{-3},10^{-2}$), Stage 2 delivers a sharp reduction in decision-time error across all tested dimensions, bringing RMSE down to the $10^{-5}$$10^{-4}$ range, while Stage 1 policies remain around $10^{-3}$. Interactive distillation consistently improves the deployable policy (Stage 1 (distill.) below Stage 1) while leaving Stage 2 essentially unchanged, confirming the intended division of labor: Stage 2 supplies a structured stationarity-correction signal, and distillation reduces the policy-class approximation/optimization gap by injecting that signal into $\pi_{\varphi}$.
Misaligned geometry. The picture becomes more heterogeneous. For small to moderate uncertainty ($s=10^{-3},10^{-2}$), Stage 2 still improves decision-time RMSE at small $d$, but its advantage shrinks with dimension and can approach the $10^{-3}$ level by $d=100$. For the largest uncertainty scale ($s=10^{-1}$), Stage 1 becomes markedly less reliable, whereas Stage 2 remains substantially better, indicating that projection can act as a stabilizing correction even when end-to-end learning is stressed. Across settings, the base and high MC budgets tend to yield similar tail-median RMSE curves, suggesting that linear-in-$d$ scaling of simulation budgets is sufficient for stable comparisons in this benchmark.
Mechanism: why misalignment can reduce projection gains. To explain when and why the projection gains shrink, we analyze Stage 2 diagnostic statistics reported in Appendix E; see Figures 36. The diagnostics indicate that the degradation under misalignment is driven primarily by increased stationarity residuals and curvature mismatch, rather than by catastrophic denominator sign failures: (i) the Stage 2 residual norm grows with dimension and becomes especially large in the hardest misaligned regime, (ii) the projection denominator magnitude stays away from zero at typical quantiles, and (iii) the bad-sign fraction remains negligible, while (iv) the effective curvature statistic $\kappa$ stays near the nominal $1/\gamma$ reference in easy regimes but can deviate substantially in the hardest misaligned/high-uncertainty setting. These patterns are consistent with the geometric explanation: when $P$ and $\Sigma$ do not commute, the inverse operations implicit in projection mix directions and can amplify Monte-Carlo errors in mixed-moment quantities (e.g., $\mathbb{E}[p_{1}\theta]$), especially as $d$ increases.
Figure 2: Pathwise sanity check at $d=100$ under common random numbers (same sampled $\theta$ and Brownian increments). Top: aligned geometry. Bottom: misaligned geometry. Each panel shows $\log X_{t}$ trajectories induced by the warm Stage 1 policy (PGDPO), the online Stage 2 P–PGDPO teacher (residual form), and the analytic $q$-reference.
Pathwise sanity check. Figure 2 complements the decision-time RMSE with a trajectory-level view under common random numbers. In the aligned case, the online Stage 2 teacher tracks the analytic $q$-reference closely along a realized path and reduces the deviation $\Delta\log X_{t}$ relative to the warm Stage 1 policy. In the misaligned case, the teacher can deviate more noticeably under the same common-noise protocol, mirroring the reduced projection advantage in the hardest regimes of Figure 1 and motivating amortization/reliability mechanisms in interactive distillation.
Overall, the benchmark highlights a separation of roles. Stage 2 projection supplies a structured stationarity-correction signal that is particularly effective under aligned uncertainty, and interactive distillation amortizes this signal into a fast deployable Stage 1 policy. Under misalignment, projection can become more sensitive as $d$ and $s$ grow, consistent with diagnostic evidence of increased residuals and curvature mismatch; nevertheless, amortization remains a robust route to improving deployable policies under fixed simulation budgets.
4.3 A strong RL baseline: PPO, and why it falls short in our benchmark
Why include PPO, and how we match the setting. Proximal Policy Optimization (PPO) is a widely used and robust model-free policy-gradient baseline for continuous control (schulman2017proximal). We include PPO to answer a concrete question: can a generic, well-tuned model-free RL method recover the decision-time $q$-optimal $\theta$-blind allocation in our high-dimensional drift-uncertainty benchmark under comparable simulation budgets? This comparison is especially informative in our static Gaussian benchmark because the target decision-time rule is structurally simple (constant and available in closed form), so performance gaps primarily reflect optimization difficulty and credit assignment rather than policy-class expressiveness.
Since classical HJB solvers and value-function-based deep PDE surrogates are not practical baselines in the high-dimensional uncertain regime targeted here (Section 2.4), PPO serves as a strong simulation-only comparator that operates on the same sampled trajectories without exploiting value-function PDE structure. For a fair comparison, PPO is trained on the same Euler simulator and time discretization as our PG–DPO pipeline, under the same deployability restriction (the policy never observes the latent $\theta$), and under the same terminal-utility objective. We also use the same action cap $u_{\max}$ (with the same dimension-scaling convention) so that exploration ranges are comparable across $d$. Implementation details are deferred to the appendix and code release.
Empirical outcome. Table LABEL:tab:crra_stage12_distill_ppo_landscape shows that PPO remains far from the analytic decision-time $q$-reference across essentially all conditions, with RMSE typically on the order of $10^{-1}$. In contrast, the Pontryagin-based pipeline attains substantially smaller errors: in aligned regimes Stage 2 projection reaches the $10^{-5}$$10^{-4}$ range for small and moderate uncertainty, while in misaligned regimes the projection advantage narrows but remains systematic. Distillation improves the deployable Stage 1 policy relative to basic PG–DPO, but does not eliminate the remaining gap to the post-hoc projection, consistent with the amortization interpretation in Section 4.2.
Why PPO underperforms in this benchmark. The gap is not evidence that PPO is intrinsically weak; rather, it reflects that our benchmark stresses regimes where a generic likelihood-ratio policy gradient is statistically disadvantaged compared to pathwise/adjoint-based updates. With terminal utility as the only reward, PPO faces a long-horizon credit-assignment problem whose gradient variance grows with both horizon and action dimension. Sampling $\theta\sim q$ further creates episode-wise heterogeneity under a single $\theta$-blind policy, inducing additional variance and potential cancellation across parameter draws. In contrast, Stage 1 exploits backpropagation through the differentiable simulator (pathwise gradients), and Stage 2 leverages the affine-in-control Pontryagin structure through a $q$-aggregated projection, replacing a noisy high-dimensional policy-gradient update by a structured stationarity correction that is tailored to the $\theta$-blind ex–ante objective.
Under matched simulation budgets in our latent-$\theta$, $\theta$-blind benchmark, a generic model-free PPO baseline does not reliably recover the decision-time $q$-optimal allocation (Table LABEL:tab:crra_stage12_distill_ppo_landscape), motivating structure-exploiting alternatives—pathwise gradients, costates, and the $q$-aggregated Pontryagin projection—as in PG–DPO and P–PGDPO.
5 Recovering Intertemporal Hedging Demand in Factor-Driven Markets
Sections 4 stressed scaling under static drift uncertainty, where the target $q$-reference is time-homogeneous and largely myopic. Here we shift the focus to an economic target: recovering the intertemporal hedging demand induced by factor-driven investment opportunities when return shocks are correlated with factor shocks (campbell2002strategic; xia2001learning).
We use the mean-reverting Gaussian premium benchmark of Section 2.3.2. Decision-time statistical uncertainty enters through the (uncertain) initial premium state $Y_{0}\sim\mathcal{N}(m_{0},P_{0})$, while a nonzero return–factor correlation $\rho$ generates hedging demand through the cross term $M_{\mathrm{cross}}(T)$ in (51). Crucially, we enforce a deployable restriction aligned with Section 3: the policy is $Y$-blind and does not observe the realized $Y_{0}$ nor the path $(Y_{t})$.
We compare: (i) Stage 1 PG–DPO (deployable end-to-end learning; Section 3.1), (ii) Stage 2 $q$-aggregated Pontryagin projection (post-hoc correction in residual form; Sections 3.2 and 3.3.1), (iii) interactive distillation (amortized projection guidance; Section 3.3.2), and (iv) a model-free PPO baseline trained under the same deployable $Y$-blind observation restriction. Performance is measured by decision-time RMSE against the analytic constant-portfolio OU reference (51) (which reduces to the independence-case benchmark when $\rho=0$).
In addition to the full allocation error, this benchmark provides a natural myopic + hedging decomposition (driven by return–factor correlation). We therefore report (a) the RMSE of the full decision-time allocation for all methods (including PPO), and (b) component-wise diagnostics for the projected (Stage 2) rules: RMSE of the hedging component (Table 1) and, in the appendix, the RMSE of the myopic component (Table 2) and the cosine similarity of the hedging direction (Table 3). Since PPO does not expose a compatible myopic/hedging decomposition for these diagnostics, we include it only in the full RMSE table.
To keep the main text focused, we include the full RMSE table (Table LABEL:tab:ou_rmse_s0_sweep_landscape) and the hedging-RMSE table (Table 1) in Section 5.2; the remaining two diagnostic tables are deferred to the appendix (Section F).
5.1 Experimental setting
$Y$-blind deployability (and what uses $Y$). Throughout this section, all reported policies are deployable and $Y$-blind: the control is a function of observable wealth and time-to-go only, and never takes the realized initial premium $Y_{0}$ nor the factor path $(Y_{t})$ as an input (including the PPO baseline). The latent premium factor is sampled and propagated only inside the simulator to generate trajectories and to form Monte Carlo averages used by the stage 2 projection (and by the teacher in distillation). Any $Y$-indexed quantities are used only for offline evaluation and diagnostics.
OU premium market with a hedging channel. We adopt the OU premium benchmark of Section 2.3.2. Let $Y_{t}\in\mathbb{R}^{m}$ be a mean-reverting premium factor and $R_{t}\in\mathbb{R}^{d}$ the risky excess returns:
$dY_{t}$ $=\kappa(\bar{y}-Y_{t})\,dt+\Xi\,dW_{t}^{Y},\hskip 18.49988ptY_{0}\sim\mathcal{N}(m_{0},P_{0}),$
$dR_{t}$ $:=\frac{dS_{t}}{S_{t}}-r\mathbf{1}\,dt=BY_{t}\,dt+\Sigma^{1/2}\,dW_{t},$
$d\langle W,W^{Y}\rangle_{t}$ $=\rho\,dt.$
A nonzero $\rho$ induces intertemporal hedging demand and enters the CRRA decision-time reference through the cross-covariance term $M_{\mathrm{cross}}(T)$ in (51). When $\rho=0$ (independent return and factor shocks), the hedging channel vanishes ($M_{\mathrm{cross}}(T)=0$) and the reference reduces to the independence-case benchmark.
Decision-time uncertainty geometry for $Y_{0}\sim\mathcal{N}(m_{0},P_{0})$. We control the magnitude of decision-time statistical uncertainty by a scalar $s_{0}>0$ and construct $P_{0}$ from an identification-motivated baseline
$\widetilde{P}_{0}:=(B^{\top}\Sigma^{-1}B)^{-1}\in\mathbb{R}^{m\times m}.$
We consider two geometries. In the aligned case, we keep the principal directions of $\widetilde{P}_{0}$ and rescale it so that the average marginal variance equals $s_{0}$:
$P_{0}^{\mathrm{aligned}}(s_{0}):=\frac{s_{0}\,m}{\mathrm{tr}(\widetilde{P}_{0})}\,\widetilde{P}_{0},$ (83)
so that $\mathrm{tr}(P_{0}^{\mathrm{aligned}})/m=s_{0}$. In the misaligned case, we preserve the eigenvalue spectrum of $\widetilde{P}_{0}$ but randomize its eigenvectors via an orthogonal rotation: letting $\widetilde{P}_{0}=U\mathrm{diag}(\lambda)U^{\top}$ be an eigen-decomposition and drawing an orthogonal matrix $R$ (e.g., Haar), we define
$P_{0}^{\mathrm{misaligned}}(s_{0}):=\frac{s_{0}\,m}{\mathrm{tr}(\widetilde{P}_{0})}\,UR\,\mathrm{diag}(\lambda)\,R^{\top}U^{\top},$ (84)
which matches the same trace normalization while rotating the uncertainty directions away from those of $\widetilde{P}_{0}$. We sweep $s_{0}\in\{10^{-3},10^{-2},10^{-1}\}$ under both aligned and misaligned $P_{0}$.
Two-stage solver, amortization, and evaluation protocol. We use the two-stage pipeline of Section 3. Stage 1 trains a deployable policy by stochastic gradient ascent using pathwise/BPTT gradients (Section 3.1). Stage 2 applies the $q$-aggregated Pontryagin projection computed under a warm-up policy (Section 3.2), implemented in the residual/control-variate form (Section 3.3.1). Interactive distillation amortizes the projected teacher into a fast deployable policy network (Section 3.3.2). As a model-free baseline, we also train a PPO policy under the same $Y$-blind observation restriction and report its decision-time full RMSE in Table LABEL:tab:ou_rmse_s0_sweep_landscape.
We sweep $d\in\{5,10,50,100\}$ (one fixed market instance per $d$), train for $5000$ epochs, and evaluate every $100$ epochs. Unless stated otherwise we set $\gamma=2$, $r=0.03$, $\kappa=1.0$, $\xi_{\mathrm{scale}}=0.25$, and $\rho=0.5$. We evaluate the decision-time action at $t=0$ and report RMSE to the analytic constant-portfolio OU reference (51).
In addition to the full allocation error, we use the natural myopic + hedging decomposition induced by the OU factor structure. We report the RMSE of the full decision-time allocation for all methods, and component-wise diagnostics for the projected (Stage 2) rules, including the RMSE of the hedging component and (in the appendix) the RMSE of the myopic component and cosine similarity of the hedging direction. To reduce noise from stochastic optimization, for each condition we summarize each metric by a tail median over the last six evaluation checkpoints.
5.2 Results: hedging-demand recovery, amortization, and robustness to decision-time uncertainty
We report decision-time RMSE at $t=0$ against the analytic OU reference (51). To reduce noise from stochastic optimization, we summarize each condition by a tail median over the last six evaluation checkpoints. Table LABEL:tab:ou_rmse_s0_sweep_landscape reports the full decision-time RMSE for all deployable objects (Stage 1 and Stage 1+Stage 2, with and without distillation), and also includes a model-free PPO baseline trained under the same $Y$-blind deployability restriction. To isolate the economic channel of interest, Table 1 reports the RMSE of the hedging component for the post-hoc projected (Stage 2) rules. Two additional diagnostics—the myopic-component RMSE and the hedging-direction cosine similarity—are deferred to the appendix (Tables 2 and 3). Since PPO does not expose a compatible myopic/hedging decomposition in our diagnostic protocol, we report it only in the full-RMSE table.
Projection and economic hedging-demand recovery. Across all $d$ and $s_{0}$, the post-hoc Pontryagin projection (Stage 1+Stage 2) substantially reduces decision-time RMSE relative to the deployable Stage 1 policy (Table LABEL:tab:ou_rmse_s0_sweep_landscape). For example, under aligned $P_{0}$ with $s_{0}=10^{-3}$ and $d=100$, Stage 1 attains $3.54\text{\times}{10}^{-03}$ whereas Stage 1+Stage 2 achieves $1.56\text{\times}{10}^{-04}$. The component-wise diagnostics indicate that the remaining discrepancy is largely driven by the hedging channel: in the same setting, the hedging RMSE is $1.55\text{\times}{10}^{-04}$ (Basic) and $1.42\text{\times}{10}^{-04}$ (Distill.) (Table 1), while the myopic RMSE is an order of magnitude smaller (Appendix Table 2). This pattern is consistent with the economic mechanism in this benchmark: once the (mostly) myopic component is captured, the dominant remaining challenge is to recover the intertemporal hedge induced by correlated return–factor shocks.
Amortization, robustness, and the PPO baseline. Interactive distillation improves the deployable Stage 1 policy relative to the basic PG–DPO run, while the most accurate object remains the post-hoc projected policy (Table LABEL:tab:ou_rmse_s0_sweep_landscape). This matches the intended division of labor in Section 3.3: Stage 2 provides a structured stationarity-correction signal through the aggregated Pontryagin projection, and distillation amortizes that correction into a single forward pass, up to policy-class approximation limits.
As the decision-time uncertainty scale $s_{0}$ increases, both the full RMSE and the hedging-component RMSE increase, with the most visible degradation at $s_{0}=10^{-1}$, especially at larger dimensions (Tables LABEL:tab:ou_rmse_s0_sweep_landscape1). Misalignment has a limited effect for small and moderate uncertainty scales, but can induce noticeable deterioration in the hardest settings, where the direction-of-hedge diagnostic can also weaken (Appendix Table 3).
Finally, the PPO baseline remains far from the analytic OU reference under the same $Y$-blind deployability restriction, with degradation that becomes especially pronounced at larger $d$ (Table LABEL:tab:ou_rmse_s0_sweep_landscape). This is consistent with PPO facing a terminal-only credit-assignment problem under latent-factor heterogeneity, in contrast to the pathwise-sensitivity and affine-in-control correction exploited by our two-stage pipeline. Since PPO does not provide a compatible myopic/hedging decomposition under our evaluation protocol, we include it only in the full-RMSE table.
$s_{0}$ Method $d=5$ $10$ $50$ $100$
Aligned $P_{0}$
$10^{-3}$ Stage 1+Stage 2 (Basic) $4.59\text{\times}{10}^{-05}$ $4.87\text{\times}{10}^{-05}$ $1.37\text{\times}{10}^{-04}$ $1.55\text{\times}{10}^{-04}$
Stage 1+Stage 2 (Distill.) $4.39\text{\times}{10}^{-05}$ $5.25\text{\times}{10}^{-05}$ $1.37\text{\times}{10}^{-04}$ $1.42\text{\times}{10}^{-04}$
$10^{-2}$ Stage 1+Stage 2 (Basic) $98\text{\times}{10}^{-05} 4.4$$47\text{\times}{10}^{-05}$ $1.44\text{\times}{10}^{-04}$ $1.74\text{\times}{10}^{-04}$
Stage 1+Stage 2 (Distill.) $4.43\text{\times}{10}^{-05}$ $4.95\text{\times}{10}^{-05}$ $1.47\text{\times}{10}^{-04}$ $1.55\text{\times}{10}^{-04}$
$10^{-1}$ Stage 1+Stage 2 (Basic) $5.27\text{\times}{10}^{-05}$ $3.99\text{\times}{10}^{-05}$ $2.60\text{\times}{10}^{-04}$ $2.95\text{\times}{10}^{-04}$
Stage 1+Stage 2 (Distill.) $4.72\text{\times}{10}^{-05}$ $5.36\text{\times}{10}^{-05}$ $2.58\text{\times}{10}^{-04}$ $2.86\text{\times}{10}^{-04}$
Misaligned $P_{0}$
$10^{-3}$ Stage 1+Stage 2 (Basic) $5.00\text{\times}{10}^{-05}$ $4.87\text{\times}{10}^{-05}$ $1.34\text{\times}{10}^{-04}$ $1.57\text{\times}{10}^{-04}$
Stage 1+Stage 2 (Distill.) $4.30\text{\times}{10}^{-05}$ $5.33\text{\times}{10}^{-05}$ $1.40\text{\times}{10}^{-04}$ $1.43\text{\times}{10}^{-04}$
$10^{-2}$ Stage 1+Stage 2 (Basic) $4.93\text{\times}{10}^{-05}$ $5.65\text{\times}{10}^{-05}$ $1.53\text{\times}{10}^{-04}$ $1.75\text{\times}{10}^{-04}$
Stage 1+Stage 2 (Distill.) $4.56\text{\times}{10}^{-05}$ $6.09\text{\times}{10}^{-05}$ $1.53\text{\times}{10}^{-04}$ $1.56\text{\times}{10}^{-04}$
$10^{-1}$ Stage 1+Stage 2 (Basic) $5.45\text{\times}{10}^{-05}$ $1.55\text{\times}{10}^{-04}$ $3.19\text{\times}{10}^{-04}$ $3.28\text{\times}{10}^{-04}$
Stage 1+Stage 2 (Distill.) $5.20\text{\times}{10}^{-05}$ $1.57\text{\times}{10}^{-04}$ $3.13\text{\times}{10}^{-04}$ $3.20\text{\times}{10}^{-04}$
Table 1: Decision-time RMSE at $t=0$ for the hedging component of the OU decision-time reference, evaluated on the post-hoc projected (Stage 2) policies (tail median over the last six evaluation checkpoints). Component-wise diagnostics are reported for Stage 2 since Stage 1 does not explicitly output a myopic/hedging decomposition.
In a factor-driven market where return–factor correlation induces intertemporal hedging demand, the proposed two-stage pipeline recovers the analytic OU decision-time reference with high accuracy: projection provides the dominant gains, and distillation improves deployable policies by amortizing projection guidance. In contrast, a model-free PPO baseline does not reliably match the analytic reference in this $Y$-blind setting.
6 Conclusion
We studied continuous-time portfolio choice in diffusion markets whose coefficients are estimated and therefore subject to statistical uncertainty (Section 2.1). We model this uncertainty by an exogenous law $q(d\theta)$ over a latent parameter $\theta$ that is drawn once at time 0 and remains fixed over the investment horizon, while the investor must deploy a single $\theta$-blind Markov feedback policy evaluated under an ex–ante CRRA objective (Remark 1, Section 2.1). This information structure shifts the relevant optimality notion from $\theta$-conditional (full-information) criticality to a $q$-aggregated Pontryagin first-order condition that is enforceable within the deployable $\theta$-blind policy class (Section 2.2, Theorem 1).
Methodologically, we extended Pontryagin–Guided Direct Policy Optimization (PG–DPO) to the latent-parameter setting by sampling $\theta$ only inside the simulator and computing exact discrete-time gradients via BPTT (Section 3.1), and we leveraged the BPTT–PMP correspondence to extract the costate objects needed for structured control updates (Theorem 2). Building on the $q$-aggregated stationarity, we proposed uncertainty-aware projected PG–DPO (P–PGDPO), which aggregates Monte Carlo Pontryagin quantities across $\theta\sim q$ and projects them onto the deployable first-order condition to obtain a single $\theta$-blind rule (Section 3.2). We established a residual-based ex–ante policy-gap bound under local stability of the aggregated projection map, with discretization and Monte Carlo errors made explicit (Theorem 3). In experiments with finite-sample uncertainty, projection improves stability and accuracy in high dimensions and exhibits a two-time-scale stabilization effect (costates versus policies), while interactive distillation amortizes the projection into a fast deployable network (Sections 4 and 5; Section 3.3).
Several extensions are natural. A first direction is to allow time-varying uncertainty descriptions $q_{t}$ (e.g., produced by an external filter) and connect the present fixed-$q$ projection to belief-aware decision rules (Remark 2, Appendix A). A second direction is to incorporate realistic frictions and constraints (transaction costs, leverage and short-sale limits) and develop certified or regularized projection steps when mixed-moment estimation becomes fragile (Section 2.4; Section 4.2; Appendix D.5). Finally, applying the framework to large cross-sectional datasets with modern estimation pipelines would further clarify the practical benefits of inference-agnostic, simulation-only optimization under parameter uncertainty (Section 1).
Acknowledgments
This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (RS-2025-00562904).
Appendix A Online uncertainty updates: Kalman–Bucy filtering and a plug-in decision-time benchmark
Purpose and scope. Sections 2.3.1 and 2.3.2 focus on decision-time benchmarks in which an uncertainty description $q$ is treated as given and the investor optimizes under the corresponding $\theta$-blind deployability constraint. In practice, however, new data arrive and the uncertainty description is updated over time by an external estimation/filtering engine, a viewpoint that aligns with learning/estimation-risk portfolio choice and Bayesian decision-time formulations (barberis2000investor; pastor2000portfolio; xia2001learning). This subsection records a simple linear–Gaussian example in which such an updated description $q_{t}$ arises endogenously via a Kalman–Bucy filter (a canonical partially observed diffusion setting; see, e.g., bensoussan1985optimal; pham2017dynamic), and then formalizes a plug-in workflow: at each decision time, treat the current uncertainty description $q_{t}$ as given and compute a decision-time optimal control under that $q_{t}$. We emphasize that solving the fully optimal partial-observation (belief-state) control problem is not the goal of this paper; rather, we view the resulting $q_{t}$ as an external input to decision-time optimization. In particular, our simulation-based Pontryagin-guided solvers developed later (Section 3) can be used as inner-loop engines that are refreshed whenever a new uncertainty description $q_{t}$ becomes available.
A linear–Gaussian hidden-premium model (OU state, observed returns). We use a stylized linear–Gaussian counterpart of the mean-reverting premium setting, but now assume the premium factor is not directly observed. Let $Y_{t}\in\mathbb{R}^{m}$ be a latent premium factor following an OU dynamics
$dY_{t}=K(\bar{y}-Y_{t})\,dt+\Xi\,dW_{t}^{Y},\hskip 18.49988ptY_{0}\sim\mathcal{N}(\hat{y}_{0},P_{0}),$ (85)
where $K\in\mathbb{R}^{m\times m}$ is stable, $\bar{y}\in\mathbb{R}^{m}$, and $\Xi\in\mathbb{R}^{m\times m}$. Risky assets satisfy
$\frac{dS_{t}}{S_{t}}=r\mathbf{1}\,dt+BY_{t}\,dt+\Sigma^{1/2}\,dW_{t},\hskip 18.49988pt\Sigma\in\mathbb{R}^{d\times d}\ \text{s.p.d.},$ (86)
with $B\in\mathbb{R}^{d\times m}$. Equivalently, the investor observes the excess-return signal
$dZ_{t}:=\frac{dS_{t}}{S_{t}}-r\mathbf{1}\,dt=BY_{t}\,dt+\Sigma^{1/2}\,dW_{t}.$ (87)
We write $\mathbb{F}^{Z}=(\mathcal{F}_{t}^{Z})_{t\in[0,T]}$ for the filtration generated by $(Z_{s})_{s\leq t}$. For clarity, we present the independent-noise case $W\perp W^{Y}$; the correlated-noise extension remains linear–Gaussian but leads to more cumbersome gain formulas.
Kalman–Bucy posterior $q_{t}=\mathcal{L}(Y_{t}\mid\mathcal{F}_{t}^{Z})$. Under (85)–(87), the conditional law of the latent factor remains Gaussian:
$q_{t}(dy):=\mathcal{L}(Y_{t}\mid\mathcal{F}_{t}^{Z})=\mathcal{N}(\hat{Y}_{t},P_{t}),$ (88)
where $(\hat{Y}_{t},P_{t})$ satisfy the Kalman–Bucy equations
$d\hat{Y}_{t}$ $=K(\bar{y}-\hat{Y}_{t})\,dt+P_{t}B^{\top}\Sigma^{-1}\Big(dZ_{t}-B\hat{Y}_{t}\,dt\Big),$ (89)
$\dot{P}_{t}$ $=KP_{t}+P_{t}K^{\top}+\Xi\Xi^{\top}-P_{t}B^{\top}\Sigma^{-1}BP_{t},\hskip 18.49988ptP_{0}\ \text{given}.$ (90)
Thus, even though the posterior $q_{t}$ is a distribution-valued object, in this affine/Gaussian regime it is fully characterized by the finite-dimensional sufficient statistics $(\hat{Y}_{t},P_{t})$, with $P_{t}$ evolving deterministically via (90); this is the prototypical setting in which belief-state control reduces to finite-dimensional sufficient statistics (bensoussan1985optimal; pham2017dynamic).
From a posterior on $Y_{t}$ to a Gaussian uncertainty description for decision-time optimization. To mirror the decision-time perspective of the OU benchmark, we consider the remaining-horizon time-averaged premium
$\bar{\theta}_{t,T}:=\frac{1}{T-t}\int_{t}^{T}BY_{s}\,ds\in\mathbb{R}^{d},\hskip 18.49988pt\tau:=T-t.$ (91)
For the OU dynamics (85), one has the decomposition
$\int_{t}^{T}Y_{s}\,ds=\tau\bar{y}+K^{-1}\big(I-e^{-K\tau}\big)(Y_{t}-\bar{y})+\int_{t}^{T}K^{-1}\big(I-e^{-K(T-u)}\big)\Xi\,dW_{u}^{Y}.$ (92)
Conditioning on $\mathcal{F}_{t}^{Z}$, the random variable $Y_{t}$ is distributed as $\mathcal{N}(\hat{Y}_{t},P_{t})$ by (88), while the future increments $(W_{u}^{Y}-W_{t}^{Y})_{u\geq t}$ are independent of $\mathcal{F}_{t}^{Z}$ in the independent-noise case. Hence $\bar{\theta}_{t,T}\mid\mathcal{F}_{t}^{Z}$ is Gaussian:
$\bar{\theta}_{t,T}\mid\mathcal{F}_{t}^{Z}\sim\mathcal{N}\!\big(m_{t,T},\,P_{t,T}\big),\hskip 18.49988ptm_{t,T}:=\frac{B\,m_{I}(t,T)}{\tau},\hskip 18.49988ptP_{t,T}:=\frac{1}{\tau^{2}}\,B\,C_{I}(t,T)\,B^{\top},$ (93)
where
$m_{I}(t,T)$ $:=\tau\bar{y}+K^{-1}\big(I-e^{-K\tau}\big)(\hat{Y}_{t}-\bar{y}),$ (94)
$C_{I}(t,T)$ $:=K^{-1}\big(I-e^{-K\tau}\big)P_{t}\big(I-e^{-K\tau}\big)^{\top}K^{-\top}+\int_{0}^{\tau}K^{-1}\big(I-e^{-Ks}\big)\Xi\Xi^{\top}\big(I-e^{-Ks}\big)^{\top}K^{-\top}\,ds.$ (95)
Equation (93) provides a concrete example of an online-updated Gaussian uncertainty description $q_{t,T}:=\mathcal{L}(\bar{\theta}_{t,T}\mid\mathcal{F}_{t}^{Z})=\mathcal{N}(m_{t,T},P_{t,T}).$
A plug-in decision-time benchmark (receding-horizon fixed-$q_{t,T}$). Given $(m_{t,T},P_{t,T})$ from (93), a simple decision-time rule is obtained by treating $q_{t,T}$ as fixed over the remaining horizon and applying the Gaussian constant-allocation benchmark of Section 2.3.1 with horizon $\tau$:
$\pi_{t}^{\mathrm{plug}}:=\Big(\gamma\,\Sigma+(\gamma-1)\tau\,P_{t,T}\Big)^{-1}m_{t,T}.$ (96)
One may interpret (96) as a receding-horizon decision-time policy driven by an externally updated uncertainty description, consistent with the general “update beliefs, then optimize” workflow used in Bayesian/learning-based portfolio choice (barberis2000investor; pastor2000portfolio; xia2001learning).
Remarks (relation to belief-aware control). The plug-in rule (96) is intentionally decision-time: it conditions on the current uncertainty description and does not attempt to optimize over how the posterior will evolve. In the present paper, we therefore treat the uncertainty law as fixed at a decision time (either as a fixed $q$ over a horizon, or as an externally updated sequence of inputs $q_{t}$ that is not controlled by the agent), mirroring the decision-time perspective common in Bayesian/learning portfolio-choice studies (barberis2000investor; pastor2000portfolio; xia2001learning). Even in linear–Gaussian regimes where the belief state is finite-dimensional, the fully optimal partial-observation portfolio problem would treat the belief state (here, $(\hat{Y}_{t},P_{t})$) as part of the controlled state and optimize the policy in that enlarged state space (bensoussan1985optimal; pham2017dynamic). Related necessary conditions under partial information can also be expressed via partial-observation maximum principles (haussmann1987maximum; li1995general; baghery2007maximum). Developing a belief-aware Pontryagin-guided policy optimizer that operates directly in $(x,y,\hat{Y},P)$-space (or its sufficient-statistic analogues) is an important direction that we defer to future work.
Appendix B Proof of Theorem 2
Theorem 2 extends the BPTT–PMP (equivalently, BPTT–BSDE) correspondence established for deterministic-parameter models in our prior work on PG–DPO (see the main BPTT–BSDE correspondence result and proof in huh2025breaking). Here the only substantive change is that the market coefficients are indexed by a random but frozen parameter $\theta\sim q$, and we need convergence statements that hold conditionally on $\theta$ and uniformly over $\theta$ in compact subsets of $\Theta$.
Important remark (what this proof does not use). This proof concerns the $\theta$-conditional Pontryagin adjoint/costate for the fixed-$\theta$ control problem induced by (53)–(54). It does not use the $\theta$-blind $q$-aggregated stationarity condition (Theorem 1 in Section 3.2). Those constructions affect only the deployable aggregation/projection target in stage 2 and are irrelevant to the BPTT–PMP convergence itself.
Notation and filtration. Fix $\theta\in\Theta$. We work conditionally on this $\theta$ and consider the augmented (simulator) filtration
$\mathbb{G}^{\theta}:=(\mathcal{G}_{t}^{\theta})_{t\in[0,T]},\hskip 18.49988pt\mathcal{G}_{t}^{\theta}:=\sigma\!\big(\theta,\{W_{s},W_{s}^{Y}:0\leq s\leq t\}\big)\ \text{(with the usual augmentation)}.$
All conditional expectations and $L^{2}$ projections below are taken with respect to $\mathcal{G}_{t_{k}}^{\theta}$. This choice matches the information set used by the simulator and by automatic differentiation/BPTT (which differentiates through the full forward recursion).
Let $\Delta t>0$, $t_{k}:=k\Delta t$, $k=0,\dots,N$, $N\Delta t=T$. For readability we suppress the policy parameters $\varphi$ and write $\pi_{k}:=\pi_{\varphi}(t_{k},X_{k}^{\theta},Y_{k}^{\theta})$, where $\pi_{\varphi}$ is $\theta$-blind in the sense of (52) but evaluated along the $\theta$-conditional trajectory.
Step 1: Conditioning on $\theta$ and uniformity of bounds. Fix a compact set $K\subset\Theta$. Assume the coefficients in (53)–(54) satisfy the usual Lipschitz and linear-growth conditions uniformly over $\theta\in K$, and that the block covariance structure of $(W,W^{Y})$ (including instantaneous correlation) is uniformly nondegenerate on $K$. Then, for each fixed $\theta\in K$, the controlled SDE system (53)–(54) is well posed and admits uniform-in-time $L^{2}$ moment bounds. Moreover, the Euler–Maruyama scheme enjoys the standard strong error bound
$\sup_{t\in[0,T]}\mathbb{E}\big[\|(X_{t}^{\pi,\theta},Y_{t}^{\theta})-(X_{t}^{\Delta t,\theta},Y_{t}^{\Delta t,\theta})\|^{2}\big]^{1/2}\;\leq\;C_{K}\,\Delta t^{1/2},$
with a constant $C_{K}$ that can be chosen independently of $\theta\in K$. These are the deterministic assumptions used in huh2025breaking, now stated uniformly on $K$.
Step 2: Discrete forward scheme and BPTT pathwise adjoints (fixed $\theta$). Under fixed $\theta$, consider the Euler scheme for (53)–(54) on the grid $(t_{k})$:
$Y_{k+1}^{\theta}=Y_{k}^{\theta}+a(Y_{k}^{\theta},\theta)\Delta t+\beta(Y_{k}^{\theta},\theta)\Delta W_{k}^{Y},$
$X_{k+1}^{\theta}=X_{k}^{\theta}+X_{k}^{\theta}\Big(r+\pi_{k}^{\top}b(Y_{k}^{\theta},\theta)\Big)\Delta t+X_{k}^{\theta}\,\pi_{k}^{\top}\sigma(Y_{k}^{\theta},\theta)\Delta W_{k},$
with terminal reward $U(X_{N}^{\theta})$. Define the discrete (pathwise) wealth costate
$p_{k}^{\mathrm{pw},\theta}:=\frac{\partial}{\partial X_{k}^{\theta}}\,U(X_{N}^{\theta}),\hskip 18.49988ptk=0,\dots,N,$
which is the same object as (59) (episode indices suppressed and dependence on $\theta$ made explicit). For the projected-control constructions we also consider the additional pathwise objects
$p_{x,k}^{\mathrm{pw},\theta}:=\frac{\partial p_{k}^{\mathrm{pw},\theta}}{\partial X_{k}^{\theta}},\hskip 18.49988ptp_{y,k}^{\mathrm{pw},\theta}:=\frac{\partial p_{k}^{\mathrm{pw},\theta}}{\partial Y_{k}^{\theta}},$
which correspond to (60). Automatic differentiation/BPTT computes $\{(p_{k}^{\mathrm{pw},\theta},p_{x,k}^{\mathrm{pw},\theta},p_{y,k}^{\mathrm{pw},\theta})\}_{k=0}^{N}$ via the backward chain rule along the discrete forward graph.
The algebraic form of the one-step backward recursion coincides with the deterministic-parameter analysis in huh2025breaking, with the replacements
$\mu\mapsto b(\cdot,\theta),\hskip 18.49988pt\sigma\mapsto\sigma(\cdot,\theta),$
and with the factor block $(Y^{\theta},\Delta W^{Y})$ handled exactly as in the wealth–factor extension therein. All one-step remainder terms are controlled by standard Taylor/Euler estimates with constants uniform in $\theta\in K$.
Step 3: One-step $L^{2}$ projection and discrete BSDE form (fixed $\theta$). Fix $\theta\in K$. As in huh2025breaking, take the conditional $L^{2}$-projection of $p_{k+1}^{\mathrm{pw},\theta}$ onto $\mathrm{span}\{1,\Delta W_{k},\Delta W_{k}^{Y}\}$ given $\mathcal{G}_{t_{k}}^{\theta}$:
$p_{k+1}^{\mathrm{pw},\theta}=\mathbb{E}\!\left[p_{k+1}^{\mathrm{pw},\theta}\mid\mathcal{G}_{t_{k}}^{\theta}\right]+z_{k}^{\theta}\Delta W_{k}+\tilde{z}_{k}^{\theta}\Delta W_{k}^{Y}+\varepsilon_{k+1}^{\theta},$
where $\varepsilon_{k+1}^{\theta}$ is orthogonal (in $L^{2}$) to $\mathrm{span}\{1,\Delta W_{k},\Delta W_{k}^{Y}\}$ conditionally on $\mathcal{G}_{t_{k}}^{\theta}$. Uniform nondegeneracy of the block covariance of $(\Delta W_{k},\Delta W_{k}^{Y})$ yields unique projection coefficients $(z_{k}^{\theta},\tilde{z}_{k}^{\theta})$.
Substituting this projection into the BPTT backward recursion from Step 2 yields a canonical discrete BSDE representation for $(p_{k}^{\mathrm{pw},\theta},z_{k}^{\theta},\tilde{z}_{k}^{\theta})$ whose drift matches the Euler discretization of the $\theta$-conditional Pontryagin adjoint BSDE associated with (53)–(54). The same argument applies to the derivatives $(p_{x,k}^{\mathrm{pw},\theta},p_{y,k}^{\mathrm{pw},\theta})$: they satisfy linearized discrete backward recursions obtained by differentiating the discrete adjoint equations, hence admit analogous discrete-BSDE representations with coefficients uniformly controlled on $K$.
Step 4: Passage to continuous time and identification with the PMP costate. For each fixed $\theta\in K$, the forward SDE and the $\theta$-conditional adjoint BSDE form a standard FBSDE with coefficients parametrized by $\theta$. Let $(p_{t}^{\theta},p_{x,t}^{\theta},p_{y,t}^{\theta})$ denote the continuous-time $\theta$-conditional Pontryagin objects under policy $\pi_{\varphi}$, so $p_{T}^{\theta}=U^{\prime}(X_{T}^{\pi,\theta})$.
Define the piecewise-constant interpolations
$p_{t}^{\Delta t,\theta}:=p_{k}^{\mathrm{pw},\theta},\qquad p_{x,t}^{\Delta t,\theta}:=p_{x,k}^{\mathrm{pw},\theta},\qquad p_{y,t}^{\Delta t,\theta}:=p_{y,k}^{\mathrm{pw},\theta},\hskip 18.49988ptt\in[t_{k},t_{k+1}).$
By the same stability and convergence arguments as in huh2025breaking (Euler convergence for the forward equation plus discrete-BSDE convergence for the backward equation), we obtain, for each fixed $\theta\in K$,
$\sup_{t\in[0,T]}\mathbb{E}\big[\,|p_{t}^{\Delta t,\theta}-p_{t}^{\theta}|^{2}\,\big]\to 0,\hskip 18.49988pt\sup_{t\in[0,T]}\mathbb{E}\big[\,\|p_{x,t}^{\Delta t,\theta}-p_{x,t}^{\theta}\|^{2}\,\big]\to 0,\hskip 18.49988pt\sup_{t\in[0,T]}\mathbb{E}\big[\,\|p_{y,t}^{\Delta t,\theta}-p_{y,t}^{\theta}\|^{2}\,\big]\to 0,$
as $\Delta t\to 0$. Because all Lipschitz, growth, ellipticity, and covariance constants were assumed uniform on $K$, the convergence constants can be chosen independently of $\theta\in K$. This yields the claimed BPTT–PMP correspondence conditionally on $\theta$ and uniformly over $\theta$ in compact subsets of $\Theta$, completing the proof. ∎
Appendix C Auxiliary results for Theorem 3
C.1 Stability of the projection map $(A,G)\mapsto-A^{-1}G$
Proposition 1 (Stability of the projection map $(A,G)\mapsto-A^{-1}G$).
Let $D$ be a measurable domain and let $\mu$ be a reference measure on $D$. Let $A,\widetilde{A}:D\to\mathbb{R}^{d\times d}$ and $G,\widetilde{G}:D\to\mathbb{R}^{d}$ be measurable. Assume:
  1. (i)
    $A(z)$ is invertible for $\mu$-a.e. $z\in D$ and $\|A^{-1}\|_{L^{\infty}(D)}\leq\kappa$ for some $\kappa>0$;
  2. (ii)
    $\|G\|_{L^{\infty}(D)}\leq M$ for some $M>0$;
  3. (iii)
    $\|\widetilde{A}-A\|_{L^{\infty}(D)}\leq(2\kappa)^{-1}$.
Define $\pi:=-A^{-1}G$ and $\widetilde{\pi}:=-\widetilde{A}^{-1}\widetilde{G}$. Then $\widetilde{A}(z)$ is invertible for $\mu$-a.e. $z\in D$ with $\|\widetilde{A}^{-1}\|_{L^{\infty}(D)}\leq 2\kappa$, and
$\|\widetilde{\pi}-\pi\|_{L^{2}(\mu)}\;\leq\;2\kappa\,\|\widetilde{G}-G\|_{L^{2}(\mu)}\;+\;2\kappa^{2}\,\big(M+\|\widetilde{G}\|_{L^{\infty}(D)}\big)\,\|\widetilde{A}-A\|_{L^{2}(\mu)}.$ (97)
Proof.
Throughout, $\|\cdot\|$ denotes the operator norm induced by the Euclidean norm. For a matrix-valued function $M:D\to\mathbb{R}^{d\times d}$, write
$\|M\|_{L^{\infty}(D)}:=\operatorname*{ess\,sup}_{z\in D}\|M(z)\|,\hskip 18.49988pt\|M\|_{L^{2}(\mu)}:=\Big(\int_{D}\|M(z)\|^{2}\,\mu(dz)\Big)^{1/2},$
and similarly for vector-valued functions.
Step 1: Invertibility and inverse bound for $\widetilde{A}$. Fix $z\in D$ such that $A(z)$ is invertible (this holds for $\mu$-a.e. $z$). Let $E(z):=\widetilde{A}(z)-A(z)$. By (i) and (iii),
$\|A(z)^{-1}E(z)\|\leq\|A(z)^{-1}\|\,\|E(z)\|\leq\|A^{-1}\|_{L^{\infty}(D)}\,\|\widetilde{A}-A\|_{L^{\infty}(D)}\leq\kappa\cdot\frac{1}{2\kappa}=\frac{1}{2}.$
Hence $I+A(z)^{-1}E(z)$ is invertible and admits the Neumann-series inverse. Therefore,
$\widetilde{A}(z)^{-1}=(A(z)+E(z))^{-1}=(I+A(z)^{-1}E(z))^{-1}A(z)^{-1},$
and
$\|\widetilde{A}(z)^{-1}\|\leq\frac{1}{1-\|A(z)^{-1}E(z)\|}\,\|A(z)^{-1}\|\leq\frac{1}{1-1/2}\,\kappa=2\kappa.$
Taking the essential supremum over $z\in D$ yields
$\|\widetilde{A}^{-1}\|_{L^{\infty}(D)}\leq 2\kappa.$
Step 2: A pointwise bound for $\widetilde{A}^{-1}-A^{-1}$. For $\mu$-a.e. $z\in D$ where both inverses exist,
$\widetilde{A}(z)^{-1}-A(z)^{-1}=\widetilde{A}(z)^{-1}\big(A(z)-\widetilde{A}(z)\big)A(z)^{-1}.$
Thus,
$\|\widetilde{A}(z)^{-1}-A(z)^{-1}\|\leq\|\widetilde{A}(z)^{-1}\|\,\|\widetilde{A}(z)-A(z)\|\,\|A(z)^{-1}\|\leq(2\kappa)\,\|\widetilde{A}(z)-A(z)\|\,\kappa=2\kappa^{2}\|\widetilde{A}(z)-A(z)\|.$
Consequently,
$\|\widetilde{A}^{-1}-A^{-1}\|_{L^{2}(\mu)}\leq 2\kappa^{2}\,\|\widetilde{A}-A\|_{L^{2}(\mu)}.$
Step 3: Control error bound. Recall $\pi=-A^{-1}G$ and $\widetilde{\pi}=-\widetilde{A}^{-1}\widetilde{G}$. Then
$\widetilde{\pi}-\pi=-\widetilde{A}^{-1}(\widetilde{G}-G)-(\widetilde{A}^{-1}-A^{-1})G.$
Taking $L^{2}(\mu)$ norms and using Hölder ($L^{\infty}\times L^{2}\to L^{2}$) gives
$\|\widetilde{\pi}-\pi\|_{L^{2}(\mu)}\leq\|\widetilde{A}^{-1}\|_{L^{\infty}(D)}\,\|\widetilde{G}-G\|_{L^{2}(\mu)}+\|\widetilde{A}^{-1}-A^{-1}\|_{L^{2}(\mu)}\,\|G\|_{L^{\infty}(D)}.$
Using $\|\widetilde{A}^{-1}\|_{L^{\infty}(D)}\leq 2\kappa$ (Step 1), $\|G\|_{L^{\infty}(D)}\leq M$ (assumption (ii)), and Step 2, we obtain
$\|\widetilde{\pi}-\pi\|_{L^{2}(\mu)}\leq 2\kappa\,\|\widetilde{G}-G\|_{L^{2}(\mu)}+2\kappa^{2}\,M\,\|\widetilde{A}-A\|_{L^{2}(\mu)}.$
Finally, since $M\leq M+\|\widetilde{G}\|_{L^{\infty}(D)}$, this implies (97). ∎
C.2 Slab-wise small-gain for the $q$-aggregated projection inputs
Time-slab decomposition.
In the portfolio problem the working domain $D$ carries a time coordinate; for concreteness, assume
$D\subset[0,T]\times\mathcal{S},\hskip 18.49988pt\mu(dt,d\xi)=dt\otimes\nu(d\xi),$
for some reference measure $\nu$ on $\mathcal{S}$. Fix a partition $0=t_{0}<t_{1}<\cdots<t_{K}=T$ with slab lengths $\tau_{k}:=t_{k}-t_{k-1}$ and define
$D_{k}:=D\cap\big([t_{k-1},t_{k}]\times\mathcal{S}\big),\hskip 18.49988pt\mu_{k}:=\mu|_{D_{k}},\hskip 18.49988pt\|f\|_{k}:=\|f\|_{L^{2}(\mu_{k})}.$
Then $\|f\|_{L^{2}(\mu)}^{2}=\sum_{k=1}^{K}\|f\|_{k}^{2}$.
Proposition 2 (Short-time (slab) Lipschitz gain).
Let $\mathcal{U}$ be a neighborhood of $\pi^{\star}$ in the deployable $\theta$-blind policy class such that for all $\pi\in\mathcal{U}$,
$\|A_{\pi}^{-1}\|_{L^{\infty}(D)}\leq\kappa,\hskip 18.49988pt\|G_{\pi}^{\mathrm{mix}}\|_{L^{\infty}(D)}\leq M_{G}.$
Assume that on each slab $D_{k}$ the $q$-aggregated projection inputs satisfy
$\|A_{\pi_{1}}-A_{\pi_{2}}\|_{k}\leq\bar{L}_{A}\,\tau_{k}^{1/2}\,\|\pi_{1}-\pi_{2}\|_{k},\hskip 18.49988pt\|G_{\pi_{1}}^{\mathrm{mix}}-G_{\pi_{2}}^{\mathrm{mix}}\|_{k}\leq\bar{L}_{G}\,\tau_{k}^{1/2}\,\|\pi_{1}-\pi_{2}\|_{k},$ (98)
for all $\pi_{1},\pi_{2}\in\mathcal{U}$ and constants $\bar{L}_{A},\bar{L}_{G}>0$ that depend only on band data. Define
$\rho(\tau):=\Big(\kappa\bar{L}_{G}+\kappa^{2}M_{G}\bar{L}_{A}\Big)\tau^{1/2}.$ (99)
Then for each slab $D_{k}$ and all $\pi_{1},\pi_{2}\in\mathcal{U}$,
$\|T(\pi_{1})-T(\pi_{2})\|_{k}\;\leq\;\rho(\tau_{k})\,\|\pi_{1}-\pi_{2}\|_{k},\hskip 18.49988ptT(\pi):=-A_{\pi}^{-1}G_{\pi}^{\mathrm{mix}}.$ (100)
In particular, if the partition is chosen with $\max_{k}\tau_{k}\leq\tau^{\star}$ for some $\tau^{\star}>0$ such that $\rho(\tau^{\star})<1$, then
$\rho_{*}:=\max_{1\leq k\leq K}\rho(\tau_{k})<1$
and $T$ is a contraction on every slab with constant at most $\rho_{*}$.
Proof.
Fix $k$ and $\pi_{1},\pi_{2}\in\mathcal{U}$. Write $A_{i}:=A_{\pi_{i}}$ and $G_{i}:=G_{\pi_{i}}^{\mathrm{mix}}$. Then
$T(\pi_{1})-T(\pi_{2})=-A_{1}^{-1}(G_{1}-G_{2})-(A_{1}^{-1}-A_{2}^{-1})G_{2},$
and
$A_{1}^{-1}-A_{2}^{-1}=A_{1}^{-1}(A_{2}-A_{1})A_{2}^{-1}.$
Using Hölder ($L^{\infty}\times L^{2}\to L^{2}$) on $D_{k}$ together with $\|A_{i}^{-1}\|_{L^{\infty}(D)}\leq\kappa$ and $\|G_{2}\|_{L^{\infty}(D)}\leq M_{G}$, we obtain
$\|T(\pi_{1})-T(\pi_{2})\|_{k}\leq\kappa\,\|G_{1}-G_{2}\|_{k}+\kappa^{2}M_{G}\,\|A_{1}-A_{2}\|_{k}.$
Applying (98) yields (100) with $\rho(\tau_{k})$ as in (99). ∎
Remark 3 (Verification of the $\tau^{1/2}$ gain and relation to prior slab analyses).
The $\tau^{1/2}$-gain in (98) is the same short-time parabolic smoothing effect used in our prior PGDPO analysis (see, e.g., huh2025breaking): one combines a Duhamel/semigroup representation of the relevant adjoint/costate objects with Young-type convolution bounds to obtain a factor $\tau^{1/2}$ on each short slab. In the present paper, the only additional bookkeeping is that $(A_{\pi},G_{\pi}^{\mathrm{mix}})$ are $q$-aggregated (in particular, linear expectations over $\theta$), which does not alter the semigroup estimates; it only changes constants through coefficient bounds uniform in $\theta$ on the compact parameter set.
C.3 Proof of Theorem 3
Setup.
Let $D$ be the working domain with reference measure $\mu$ and slab decomposition $\{D_{k},\mu_{k},\|\cdot\|_{k}\}_{k=1}^{K}$ as in Appendix C.2. For a deployable $\theta$-blind policy $\pi$, write $(A_{\pi},G_{\pi}^{\mathrm{mix}})$ for the $q$-aggregated projection inputs corresponding to the mixed-moment aggregation in (65)–(66). Define the projection map
$T(\pi)(z):=-A_{\pi}(z)^{-1}G_{\pi}^{\mathrm{mix}}(z),\hskip 18.49988ptz\in D,$
whenever $A_{\pi}(z)$ is invertible.
Let $\pi^{\mathrm{warm}}$ be the warm-up policy and set
$A_{\mathrm{warm}}:=A_{\pi^{\mathrm{warm}}},\hskip 18.49988ptG_{\mathrm{warm}}:=G_{\pi^{\mathrm{warm}}}^{\mathrm{mix}},\hskip 18.49988pt\pi_{\mathrm{proj}}:=T(\pi^{\mathrm{warm}})=-A_{\mathrm{warm}}^{-1}G_{\mathrm{warm}}.$
Let $\widehat{A}_{t}$ and $\widehat{G}_{t}^{\mathrm{mix}}$ be the BPTT/Monte Carlo estimators constructed under $\pi^{\mathrm{warm}}$, and denote
$\widehat{\pi}^{\mathrm{agg,mix}}:=-\widehat{A}_{t}^{-1}\widehat{G}_{t}^{\mathrm{mix}}\hskip 18.49988pt\text{on }D.$
Proof.
Step 0 (Fixed-point form of the deployable optimum). Let $\pi^{\star}:=\pi^{\star,\mathrm{blind}}$ be a locally optimal interior deployable $\theta$-blind policy for the fixed-$q$ ex–ante problem. By the $q$-aggregated stationarity (Theorem 1), $\pi^{\star}$ satisfies
$A_{\pi^{\star}}(z)\,\pi^{\star}(z)=-G_{\pi^{\star}}^{\mathrm{mix}}(z)\qquad\text{for }\mu\text{-a.e.\ }z\in D,$
hence (under invertibility on $D$) it is a fixed point of $T$:
$\pi^{\star}(z)=T(\pi^{\star})(z)=-A_{\pi^{\star}}(z)^{-1}G_{\pi^{\star}}^{\mathrm{mix}}(z),\hskip 18.49988pt\mu\text{-a.e.\ }z\in D.$
Step 1 (Triangle decomposition). Add and subtract $\pi_{\mathrm{proj}}$:
$\|\widehat{\pi}^{\mathrm{agg,mix}}-\pi^{\star}\|_{L^{2}(\mu)}\leq\|\widehat{\pi}^{\mathrm{agg,mix}}-\pi_{\mathrm{proj}}\|_{L^{2}(\mu)}+\|\pi_{\mathrm{proj}}-\pi^{\star}\|_{L^{2}(\mu)}.$ (101)
Step 2 (Estimation error via Proposition 1). Apply Proposition 1 with
$(A,G)=(A_{\mathrm{warm}},G_{\mathrm{warm}}),\hskip 18.49988pt(\widetilde{A},\widetilde{G})=(\widehat{A}_{t},\widehat{G}_{t}^{\mathrm{mix}}).$
Assume the perturbative regime on $D$:
$\|A_{\mathrm{warm}}^{-1}\|_{L^{\infty}(D)}\leq\kappa,\qquad\|G_{\mathrm{warm}}\|_{L^{\infty}(D)}\leq M_{G},\qquad\|\widehat{A}_{t}-A_{\mathrm{warm}}\|_{L^{\infty}(D)}\leq(2\kappa)^{-1},\qquad\|\widehat{G}_{t}^{\mathrm{mix}}\|_{L^{\infty}(D)}\leq M_{G}.$ (102)
Then (97) yields
$\|\widehat{\pi}^{\mathrm{agg,mix}}-\pi_{\mathrm{proj}}\|_{L^{2}(\mu)}$ $\leq 2\kappa\,\|\widehat{G}_{t}^{\mathrm{mix}}-G_{\mathrm{warm}}\|_{L^{2}(\mu)}+4\kappa^{2}M_{G}\,\|\widehat{A}_{t}-A_{\mathrm{warm}}\|_{L^{2}(\mu)}.$ (103)
By the definition of $\delta_{\mathrm{BPTT}}(\Delta t,M_{\mathrm{MC}},M_{\theta})$ in Theorem 3,
$\|\widehat{A}_{t}-A_{\mathrm{warm}}\|_{L^{2}(\mu)}+\|\widehat{G}_{t}^{\mathrm{mix}}-G_{\mathrm{warm}}\|_{L^{2}(\mu)}\leq\delta_{\mathrm{BPTT}}(\Delta t,M_{\mathrm{MC}},M_{\theta}),$
hence
$\|\widehat{\pi}^{\mathrm{agg,mix}}-\pi_{\mathrm{proj}}\|_{L^{2}(\mu)}\leq C_{2}\,\delta_{\mathrm{BPTT}}(\Delta t,M_{\mathrm{MC}},M_{\theta}),\hskip 18.49988ptC_{2}:=2\kappa+4\kappa^{2}M_{G}.$ (104)
Step 3 (Slab-wise warm-up bias bound). Assume $\pi^{\mathrm{warm}},\pi^{\star}\in\mathcal{U}$ and the slab-wise contraction $\|T(\pi_{1})-T(\pi_{2})\|_{k}\leq\rho(\tau_{k})\|\pi_{1}-\pi_{2}\|_{k}$ from Proposition 2. Let $\rho_{*}:=\max_{k}\rho(\tau_{k})<1$. Since $\pi_{\mathrm{proj}}=T(\pi^{\mathrm{warm}})$ and $\pi^{\star}=T(\pi^{\star})$, for each slab $D_{k}$ we have
$\|\pi_{\mathrm{proj}}-\pi^{\star}\|_{k}=\|T(\pi^{\mathrm{warm}})-T(\pi^{\star})\|_{k}\leq\rho(\tau_{k})\,\|\pi^{\mathrm{warm}}-\pi^{\star}\|_{k}\leq\rho_{*}\,\|\pi^{\mathrm{warm}}-\pi^{\star}\|_{k}.$ (105)
Step 4 (Residual identity and slab-wise closure). Define the warm-up aggregated stationarity residual (mixed-moment) on $D$ by
$r_{\mathrm{FOC,mix}}^{\mathrm{warm}}(z):=A_{\mathrm{warm}}(z)\,\pi^{\mathrm{warm}}(z)+G_{\mathrm{warm}}(z),\hskip 18.49988pt\varepsilon_{\mathrm{warm}}^{\mathrm{mix}}:=\|r_{\mathrm{FOC,mix}}^{\mathrm{warm}}\|_{L^{2}(\mu)}.$
Also define the slab-wise residual sizes
$\varepsilon_{\mathrm{warm},k}^{\mathrm{mix}}:=\|r_{\mathrm{FOC,mix}}^{\mathrm{warm}}\|_{k},\hskip 18.49988pt\text{so that }\ (\varepsilon_{\mathrm{warm}}^{\mathrm{mix}})^{2}=\sum_{k=1}^{K}(\varepsilon_{\mathrm{warm},k}^{\mathrm{mix}})^{2}.$
By construction of $\pi_{\mathrm{proj}}$,
$\pi^{\mathrm{warm}}-\pi_{\mathrm{proj}}=A_{\mathrm{warm}}^{-1}r_{\mathrm{FOC,mix}}^{\mathrm{warm}},$
hence on each slab
$\|\pi^{\mathrm{warm}}-\pi_{\mathrm{proj}}\|_{k}\leq\kappa\,\varepsilon_{\mathrm{warm},k}^{\mathrm{mix}}.$ (106)
Now combine the triangle inequality on each slab with (105):
$\|\pi^{\mathrm{warm}}-\pi^{\star}\|_{k}\leq\|\pi^{\mathrm{warm}}-\pi_{\mathrm{proj}}\|_{k}+\|\pi_{\mathrm{proj}}-\pi^{\star}\|_{k}\leq\kappa\,\varepsilon_{\mathrm{warm},k}^{\mathrm{mix}}+\rho_{*}\,\|\pi^{\mathrm{warm}}-\pi^{\star}\|_{k}.$
Since $\rho_{*}<1$, we close slab-wise:
$\|\pi^{\mathrm{warm}}-\pi^{\star}\|_{k}\leq\frac{\kappa}{1-\rho_{*}}\,\varepsilon_{\mathrm{warm},k}^{\mathrm{mix}}.$ (107)
Plugging into (105) gives
$\|\pi_{\mathrm{proj}}-\pi^{\star}\|_{k}\leq\frac{\rho_{*}\kappa}{1-\rho_{*}}\,\varepsilon_{\mathrm{warm},k}^{\mathrm{mix}}.$ (108)
Summing over slabs yields the global bias bound
$\|\pi_{\mathrm{proj}}-\pi^{\star}\|_{L^{2}(\mu)}\leq\frac{\rho_{*}\kappa}{1-\rho_{*}}\,\varepsilon_{\mathrm{warm}}^{\mathrm{mix}}.$ (109)
Step 5 (Finish). Combine (101), (104), and (109) to obtain
$\|\widehat{\pi}^{\mathrm{agg,mix}}-\pi^{\star}\|_{L^{2}(\mu)}\leq\frac{\rho_{*}\kappa}{1-\rho_{*}}\,\varepsilon_{\mathrm{warm}}^{\mathrm{mix}}+C_{2}\,\delta_{\mathrm{BPTT}}(\Delta t,M_{\mathrm{MC}},M_{\theta}),$
which is the slab-wise version of (71) (with $\rho_{*}$ in place of a global $C_{1}$). ∎
Remark 4 (Relation to prior PGDPO slab analyses).
The closure step above uses the same slab-wise small-gain philosophy as in huh2025breaking: short-time parabolic smoothing yields a contraction on each time slab, and the global bound follows by concatenation. The key difference here is that the contraction is applied to the $q$-aggregated projection map $T(\pi)=-A_{\pi}^{-1}G_{\pi}^{\mathrm{mix}}$, hence the additional use of the algebraic projection stability (Proposition 1) for the estimator $\widehat{\pi}^{\mathrm{agg,mix}}$.
Appendix D Implementation details for Section 3
This appendix provides reproducible step-by-step templates for the methods in Section 3. The high-level pipeline is summarized in Figure LABEL:fig:sec3-pipeline-sideways. For compactness we present one template per subsection of Section 3.
D.1 Stage 1 (PG–DPO) template for Section 3.1
Stage 1 performs stochastic gradient ascent on the fixed-$q$ ex–ante objective (56), with latent $\theta\sim q$ sampled inside the simulator while the policy remains $\theta$-blind.
Inputs.
Policy parameters $\varphi$; sampler $\nu$ over initial states; prior $q$; time grid $(N,\Delta t)$; batch size $M$; optimizer and step size $\alpha$.
Template (one training iteration).
  1. 1.
    Sample initial states. Draw a mini-batch $\{z_{0}^{(i)}=(t_{0}^{(i)},x_{0}^{(i)},y_{0}^{(i)})\}_{i=1}^{M}\sim\nu$.
  2. 2.
    Sample latent environment parameter. Sample $\theta\sim q$ inside the simulator (unseen by $\pi_{\varphi}$). (Variant: sample $\theta^{(i)}\sim q$ independently per episode; both are unbiased for $\nabla_{\varphi}J(\varphi)$.)
  3. 3.
    Simulate Euler rollouts. For each episode $i$, simulate the Euler scheme in (53)–(54) under the $\theta$-blind policy $\pi_{\varphi}$ and collect terminal utilities $\{U(X_{T}^{(i)})\}_{i=1}^{M}$.
  4. 4.
    Backpropagation through time (BPTT). Compute the Monte Carlo gradient estimator
    $\widehat{g}\;\leftarrow\;\frac{1}{M}\sum_{i=1}^{M}\nabla_{\varphi}U(X_{T}^{(i)}).$
  5. 5.
[…]
  6. 6.
    Checkpoint. Periodically save a warm-up checkpoint $\varphi^{\mathrm{warm}}$ for stage 2 projection.
D.2 Stage 2 (P–PGDPO projection; mixed-moment $q$-aggregation) template for Section 3.2
Stage 2 is a post-processing map: given a warm-up $\theta$-blind policy $\pi_{\varphi^{\mathrm{warm}}}$, it estimates Pontryagin sensitivity objects by Monte Carlo and constructs a deployable projected control on a working-domain sample $z\sim\mu$.
The aggregation used here matches the mixed-moment $q$-aggregation in (65)–(66), yielding the projected control (67).
Inputs.
Warm-up policy $\pi_{\varphi^{\mathrm{warm}}}$; working-domain sampler $\mu$ on $D$; budgets $(M_{z},M_{\theta},M_{\mathrm{MC}})$.
Template (constructing projection targets on a batch of query states).
  1. 1.
    Sample working-domain query states. Draw $\{z_{j}=(t_{j},x_{j},y_{j})\}_{j=1}^{M_{z}}\sim\mu$.
  2. 2.
    For each query state $z_{j}$, sample latent parameters. Sample $\{\theta_{\ell}\}_{\ell=1}^{M_{\theta}}\sim q$.
  3. 3.
    For each frozen $\theta_{\ell}$, estimate costates at $z_{j}$. For each $\ell=1,\dots,M_{\theta}$:
    1. (a)
      Simulate $M_{\mathrm{MC}}$ trajectories from $z_{j}$ under $\pi_{\varphi^{\mathrm{warm}}}$ with frozen $\theta_{\ell}$.
    2. (b)
      Compute pathwise sensitivities by autodiff/BPTT and average as in (62) to obtain $\widehat{p}_{t}^{\theta_{\ell}}(z_{j})$, $\widehat{p}_{x,t}^{\theta_{\ell}}(z_{j})$, $\widehat{p}_{y,t}^{\theta_{\ell}}(z_{j})$.
    3. (c)
      Form the $\theta$-conditional inputs (cf. (63)–(64)):
      $\widehat{A}_{t}^{\theta_{\ell}}(z_{j})\leftarrow x_{j}\,\widehat{p}_{x,t}^{\theta_{\ell}}(z_{j})\,\Sigma(y_{j},\theta_{\ell}),\hskip 18.49988pt\widehat{G}_{t}^{\theta_{\ell}}(z_{j})\leftarrow\widehat{p}_{t}^{\theta_{\ell}}(z_{j})\,b(y_{j},\theta_{\ell})+\Sigma_{SY}(y_{j},\theta_{\ell})\,\widehat{p}_{y,t}^{\theta_{\ell}}(z_{j}).$
  4. 4.
    Aggregate across $\theta\sim q$ (mixed-moment). Compute
    $\widehat{A}_{t}(z_{j})\;\leftarrow\;\frac{1}{M_{\theta}}\sum_{\ell=1}^{M_{\theta}}\widehat{A}_{t}^{\theta_{\ell}}(z_{j}),\hskip 18.49988pt\widehat{G}_{t}^{\mathrm{mix}}(z_{j})\;\leftarrow\;\frac{1}{M_{\theta}}\sum_{\ell=1}^{M_{\theta}}\widehat{G}_{t}^{\theta_{\ell}}(z_{j}),$
    consistent with (65)–(66).
  5. 5.
    Solve the projection (mixed-moment aggregation). Whenever $\widehat{A}_{t}(z_{j})$ is invertible and the solve is numerically stable, compute the deployable projected control
    $\widehat{\pi}^{\mathrm{agg,mix}}(z_{j})\;\leftarrow\;-\big(\widehat{A}_{t}(z_{j})\big)^{-1}\widehat{G}_{t}^{\mathrm{mix}}(z_{j}),$
    which matches (67).
D.3 Coupling I: residual/control-variate projection (Section 3.3.1)
This subsection records a variance-reduced implementation of the stage 2 map using the residual identity (74). The residual form is applied around the warm-up policy and uses the mixed-moment aggregated inputs $(\widehat{A}_{t},\widehat{G}_{t}^{\mathrm{mix}})$.
Inputs.
Warm-up policy $\pi_{\varphi^{\mathrm{warm}}}$; query state(s) $z=(t,x,y)\sim\mu$; and stage 2 projection ingredients $(\widehat{A}_{t}(z),\widehat{G}_{t}^{\mathrm{mix}}(z))$ constructed as in Section D.2.
Template (statewise residual projection; mixed-moment aggregation).
  1. 1.
    Evaluate warm-up control. Compute $\pi_{\varphi^{\mathrm{warm}}}(z)$.
  2. 2.
    Form the aggregated residual. Compute
    $\widehat{r}_{\mathrm{FOC}}(z)\;\leftarrow\;\widehat{A}_{t}(z)\,\pi_{\varphi^{\mathrm{warm}}}(z)+\widehat{G}_{t}^{\mathrm{mix}}(z).$
  3. 3.
    Apply the residual correction. Compute
    $\widehat{\pi}^{\mathrm{agg,mix}}(z)\;\leftarrow\;\pi_{\varphi^{\mathrm{warm}}}(z)-\big(\widehat{A}_{t}(z)\big)^{-1}\widehat{r}_{\mathrm{FOC}}(z).$
D.4 Coupling II: interactive distillation (Section 3.3.2)
This subsection records an implementation template for interactive distillation: the projected output from stage 2 is used as a teacher signal during stage 1 training via the mixed objective (76). The teacher is built from the mixed-moment projected rule (possibly evaluated in residual form for variance reduction).
Inputs.
Student parameters $\varphi$; teacher refresh interval $K$; distillation schedule $\lambda(n)$; working-domain sampler $\mu$.
Template (training loop with intermittent teacher refresh).
  1. 1.
    Initialize. Set $\varphi^{-}\leftarrow\varphi$ and initialize an empty teacher buffer $\mathcal{B}\leftarrow\emptyset$.
  2. 2.
    Repeat for iterations $n=1,2,\dots$:
    1. (a)
      Stage 1 update (PG–DPO step). Perform one PG–DPO update step on $J(\varphi)$ as in Section D.1.
    2. (b)
      Teacher refresh (every $K$ steps). If $n\bmod K=0$:
      1. i.
        Set $\varphi^{-}\leftarrow\varphi$ (lagged copy).
      2. ii.
        Sample working-domain states $\{z_{j}\}_{j=1}^{M_{z}}\sim\mu$.
      3. iii.
        For each $z_{j}$, run stage 2 under $\pi_{\varphi^{-}}$ (mixed-moment aggregation) to compute a projected teacher $\widehat{\pi}^{\mathrm{agg,mix}}_{\varphi^{-}}(z_{j})$. (In practice we compute it in residual form around $\pi_{\varphi^{-}}$ as in (75).)
      4. iv.
        Optionally filter states using diagnostics (Section D.5) and update the buffer:
        $\mathcal{B}\ \leftarrow\ \{(z_{j},\widehat{\pi}^{\mathrm{agg,mix}}_{\varphi^{-}}(z_{j}))\}_{j=1}^{M_{z}}\ \ \text{(after filtering)}.$
    3. (c)
      Distillation step (when enabled). If $\lambda(n)>0$ and $\mathcal{B}\neq\emptyset$:
      1. i.
        Sample $(z,\pi^{\mathrm{teach}})$ from $\mathcal{B}$.
      2. ii.
        Apply a gradient step to minimize the proximity term $\|\pi_{\varphi}(z)-\mathrm{stopgrad}(\pi^{\mathrm{teach}})\|^{2}$ with coefficient $\lambda(n)$, consistent with (76).
D.5 Engineering notes and stabilizers
This subsection collects practical stabilizers that we found helpful for reliable training and projection in high dimensions.
  • Antithetic sampling for $\theta$. When $q$ is symmetric (e.g. Gaussian in a latent normal parameterization), sample $\theta$ in antithetic pairs by drawing $z\sim\mathcal{N}(0,I)$ and using (z,z)(z,-z) to construct $(\theta^{+},\theta^{-})$. This reduces the variance of $q$-averaged quantities and typically improves the stability of stage 2 diagnostics on the working domain.
  • Blockwise Monte Carlo and robust aggregation. To control rare-tail domination, split Monte Carlo replications into $B$ blocks and compute blockwise averages of costate-driven ingredients (e.g. $\widehat{A}_{t}^{\theta}(z)$ and $\widehat{G}_{t}^{\theta}(z)$). Aggregate across blocks using a robust statistic such as the median or median-of-means, which makes the projection less sensitive to outlier trajectories.
  • Curvature/denominator stability checks. Because the projection map $(A,G)\mapsto-A^{-1}G$ can be sensitive to near-singularity of $A$, monitor the conditioning of $\widehat{A}_{t}$ (or failure rates of the linear solve). When diagnostics indicate ill-conditioning, skip projection-guided updates at that state or increase Monte Carlo budgets locally.
  • Residual magnitude as a reliability diagnostic. For the residual form, compute $\widehat{r}_{\mathrm{FOC}}(z)=\widehat{A}_{t}(z)\pi_{\varphi^{\mathrm{warm}}}(z)+\widehat{G}_{t}^{\mathrm{mix}}(z)$. Small $\|\widehat{r}_{\mathrm{FOC}}(z)\|$ indicates approximate satisfaction of the mixed-moment aggregated first-order condition at $z$ and empirically correlates with more reliable teacher targets.
  • Diagnostics-based teacher selection on the working domain. Rather than applying distillation on all sampled $\{z_{j}\}\sim\mu$, keep only states that pass a reliability predicate. In practice, filter using residual-magnitude thresholds together with stable linear-solve diagnostics to prevent a small subset of pathological states from contaminating the teacher buffer.
  • $\lambda$ schedule and safeguards. Use a warm-up period with $\lambda=0$ (pure PG–DPO) and increase $\lambda$ only after stage 2 diagnostics on the working domain are stable. To prevent the teacher term from dominating the ex–ante objective, cap the effective coefficient via
    $\lambda_{\mathrm{eff}}:=\min\Big\{\lambda,\;c\,\frac{|L_{\mathrm{main}}|}{L_{\mathrm{distill}}+\varepsilon}\Big\},$
    with $c\in(0,1)$ and $\varepsilon>0$.
  • Initialization and scale control in high dimensions. To avoid early-time numerical blow-ups (often through quadratic variation terms of the form $\pi^{\top}\Sigma\pi$), initialize the policy output near zero and/or scale the output by $d^{-1/2}$. As a last-resort safety net, a mild log-wealth clamp can prevent overflow, but it should be used conservatively and monitored, since frequent clamping may distort higher-order sensitivities.
Appendix E Stage 2 projection diagnostics
We report Stage 2 diagnostic statistics as a visual supplement to Section 4.2. Each figure summarizes the same tail-median protocol and layout; see captions for definitions and interpretation.
Figure 3: Stage 2 stationarity residual (q50). All panels report tail medians over epochs 9500–10000 (final six evaluation snapshots). Layout matches Figure 1: rows correspond to $s\in\{10^{-3},10^{-2},10^{-1}\}$ and columns correspond to aligned vs. misaligned uncertainty. Solid vs. dashed lines are MC base ($100\cdot d$) vs. high ($400\cdot d$). We plot the median (q50) of the estimated Hamiltonian first-order condition residual norm at the query states. Larger residual indicates the warm policy is farther from stationarity, implying a larger correction is required in the residual-form projection. Growth of this residual with $d$ (especially under misalignment) supports the mechanism that projection becomes more sensitive in high dimension due to larger correction magnitudes and amplified mixed-moment noise.
Figure 4: Stage 2 denominator magnitude (q50). All panels report tail medians over epochs 9500–10000 (final six evaluation snapshots). Layout matches Figure 1: rows correspond to $s\in\{10^{-3},10^{-2},10^{-1}\}$ and columns correspond to aligned vs. misaligned uncertainty. Solid vs. dashed lines are MC base ($100\cdot d$) vs. high ($400\cdot d$). We plot a typical (q50) magnitude of the projection denominator/curvature term used in the residual-form update. Values bounded away from zero indicate that projection is not operating in a near-singular regime at typical quantiles. This helps rule out “catastrophic inversion” as the primary driver of degradation; instead, residual growth and curvature mismatch (Fig. 5) provide a more consistent explanation in misaligned/high-$d$ regimes.
Figure 5: Stage 2 curvature-consistency statistic $\kappa$. All panels report tail medians over epochs 9500–10000 (final six evaluation snapshots). Layout matches Figure 1: rows correspond to $s\in\{10^{-3},10^{-2},10^{-1}\}$ and columns correspond to aligned vs. misaligned uncertainty. Solid vs. dashed lines are MC base ($100\cdot d$) vs. high ($400\cdot d$). We report the stabilized median-after-floor statistic $\kappa$ and compare it to the nominal reference $1/\gamma$ (horizontal dotted line). For CRRA, costate ratios imply a characteristic curvature scale; sustained deviations of $\kappa$ from $1/\gamma$ indicate costate inconsistency and/or bias in mixed-moment estimation, and are most visible in the hardest misaligned/high-uncertainty regime.
Figure 6: Stage 2 bad-sign fraction. All panels report tail medians over epochs 9500–10000 (final six evaluation snapshots). Layout matches Figure 1: rows correspond to $s\in\{10^{-3},10^{-2},10^{-1}\}$ and columns correspond to aligned vs. misaligned uncertainty. Solid vs. dashed lines are MC base ($100\cdot d$) vs. high ($400\cdot d$). We plot the fraction of samples in which the estimated curvature/denominator violates the expected sign condition (loss of local concavity on the sampled batch). Near-zero bad-sign fractions across most regimes suggest that the projection typically operates in a locally well-behaved region and that failures are not dominated by sign flips, supporting the main-text conclusion that misalignment primarily increases residual/costate mismatch rather than inducing widespread concavity violations.
Appendix F Supplementary decomposition diagnostics for Section 5
Tables 23 report Stage 2 decomposition diagnostics at $t=0$.
$s_{0}$ Method $d=5$ $10$ $50$ $100$
Aligned $P_{0}$
$10^{-3}$ Stage 1+Stage 2 (Basic) $7.17\text{\times}{10}^{-06}$ $1.51\text{\times}{10}^{-05}$ $1.60\text{\times}{10}^{-05}$ $1.23\text{\times}{10}^{-05}$
Stage 1+Stage 2 (Distill.) $5.20\text{\times}{10}^{-06}$ $9.81\text{\times}{10}^{-06}$ $1.64\text{\times}{10}^{-05}$ $1.62\text{\times}{10}^{-05}$
$10^{-2}$ Stage 1+Stage 2 (Basic) $6.21\text{\times}{10}^{-06}$ $1.38\text{\times}{10}^{-05}$ $1.63\text{\times}{10}^{-05}$ $1.42\text{\times}{10}^{-05}$
Stage 1+Stage 2 (Distill.) $7.13\text{\times}{10}^{-06}$ $7.13\text{\times}{10}^{-06}$ $1.62\text{\times}{10}^{-05}$ $1.76\text{\times}{10}^{-05}$
$10^{-1}$ Stage 1+Stage 2 (Basic) $8.18\text{\times}{10}^{-06}$ $1.41\text{\times}{10}^{-05}$ $3.82\text{\times}{10}^{-05}$ $2.42\text{\times}{10}^{-05}$
Stage 1+Stage 2 (Distill.) $6.72\text{\times}{10}^{-06}$ $9.21\text{\times}{10}^{-06}$ $3.67\text{\times}{10}^{-05}$ $3.16\text{\times}{10}^{-05}$
Misaligned $P_{0}$
$10^{-3}$ Stage 1+Stage 2 (Basic) $1.10\text{\times}{10}^{-05}$ $1.41\text{\times}{10}^{-05}$ $70\text{\times}{10}^{-05} 1.1$$18\text{\times}{10}^{-05}$
Stage 1+Stage 2 (Distill.) $7.71\text{\times}{10}^{-06}$ $5.94\text{\times}{10}^{-06}$ $1.84\text{\times}{10}^{-05}$ $1.62\text{\times}{10}^{-05}$
$10^{-2}$ Stage 1+Stage 2 (Basic) $7.90\text{\times}{10}^{-06}$ $2.02\text{\times}{10}^{-05}$ $1.46\text{\times}{10}^{-05}$ $1.20\text{\times}{10}^{-05}$
Stage 1+Stage 2 (Distill.) $6.00\text{\times}{10}^{-06}$ $1.71\text{\times}{10}^{-05}$ $2.21\text{\times}{10}^{-05}$ $1.43\text{\times}{10}^{-05}$
$10^{-1}$ Stage 1+Stage 2 (Basic) $1.24\text{\times}{10}^{-05}$ $1.93\text{\times}{10}^{-04}$ $5.77\text{\times}{10}^{-05}$ $3.14\text{\times}{10}^{-05}$
Stage 1+Stage 2 (Distill.) $1.20\text{\times}{10}^{-05}$ $1.90\text{\times}{10}^{-04}$ $7.40\text{\times}{10}^{-05}$ $2.47\text{\times}{10}^{-05}$
Table 2: Myopic-component RMSE at $t=0$ (tail medians).
$s_{0}$ Method $d=5$ $10$ $50$ $100$
Aligned $P_{0}$
$10^{-3}$ Stage 1+Stage 2 (Basic) $0.994$ $0.988$ $0.991$ $0.990$
Stage 1+Stage 2 (Distill.) $0.995$ $0.986$ $0.990$ $0.987$
$10^{-2}$ Stage 1+Stage 2 (Basic) $0.993$ $0.989$ $0.992$ $0.988$
Stage 1+Stage 2 (Distill.) $0.992$ $0.994$ $0.990$ $0.987$
$10^{-1}$ Stage 1+Stage 2 (Basic) $0.988$ $0.990$ $0.936$ $0.932$
Stage 1+Stage 2 (Distill.) $0.996$ $0.990$ $0.949$ $0.922$
Misaligned $P_{0}$
$10^{-3}$ Stage 1+Stage 2 (Basic) $0.988$ $0.988$ $0.993$ $0.990$
Stage 1+Stage 2 (Distill.) $0.994$ $0.995$ $0.990$ $0.987$
$10^{-2}$ Stage 1+Stage 2 (Basic) $0.994$ $0.976$ $0.992$ $0.988$
Stage 1+Stage 2 (Distill.) $0.994$ $0.980$ $0.988$ $0.989$
$10^{-1}$ Stage 1+Stage 2 (Basic) $0.990$ $0.005$ $0.668$ $0.851$
Stage 1+Stage 2 (Distill.) $0.992$ 0.009-0.009 $0.642$ $0.871$
Table 3: Hedging-direction cosine similarity at $t=0$ (tail medians). Higher is better; negative indicates direction reversal.